### v2

# The Complexity of Minimum-Envy House Allocation Over Graphs

Abstract
In this paper, we study a generalization of the House Allocation problem. In our problem, agents are represented by vertices of a graph $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$, and each agent $a\in\mathcal{A}$ is associated with a set of preferred houses $\mathcal{P}_{a}\subseteq\mathcal{H}$, where $\mathcal{A}$ is the set of agents and $\mathcal{H}$ is the set of houses. A house allocation is an injective function $\phi:\mathcal{A}\rightarrow\mathcal{H}$, and an agent $a$ envies a neighbour $a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)$ under $\phi$ if $\phi(a)\notin\mathcal{P}_{a}$ and $\phi(a^{\prime})\in\mathcal{P}_{a}$. We study two natural objectives: the first problem called Optimal House Allocation of Agent Network, aims to compute an allocation that minimizes the number of envious agents; the second problem called Optimally Happy House Allocation of Agent Network aims to maximize, among all minimum-envy allocations, the number of agents who are assigned a house they prefer. These two objectives capture complementary notions of fairness and individual satisfaction.
Our main results provide a comprehensive algorithmic and complexity-theoretic understanding of these objectives. First, we design polynomial time algorithms for both problems for the variant when each agent prefers exactly one house. On the other hand, when the list of preferred houses for each agent has size at most $2$ then we show that both problems are $\mathsf{NP}$-hard even when the agent graph $\mathcal{G}_{\mathcal{A}}$ is a complete bipartite graph. We also show that both problems are $\mathsf{NP}$-hard even when the number $|\mathcal{H}|$ of houses is equal to the number $|\mathcal{A}|$ of agents. This is in contrast to the classical House Allocation problem, where the problem is polynomial time solvable when $|\mathcal{H}|=|\mathcal{A}|$. The two problems are also $\mathsf{NP}$-hard when the agent graph has a small vertex cover. On the positive side, we design exact algorithms that exploit certain structural properties of $\mathcal{G}_{\mathcal{A}}$ such as sparsity, existence of balanced separators or existence of small-sized vertex covers, and perform better than the naive brute-force algorithm.
keywords:
fair allocation, balanced separators, envy-freeness
category:
[…]
1 Introduction
The challenge of allocating limited resources in a fair and efficient manner arises in a wide variety of real-world contexts, from assigning dorm rooms to students, to distributing public housing, to allocating virtual machines in cloud systems. A particularly well-studied problem in this area is the House Allocation problem, in which we are given a set $\mathcal{A}$ of agents and a set $\mathcal{H}$ of houses. An allocation of houses to agents is simply an injective function from agents to houses. We are also given for each agent $a$ a list $\mathcal{P}_{a}$ of preferred houses. The agent is satisfied if they are allocated a house from $\mathcal{P}_{a}$. On the other hand, if agent $a$ is not allocated a house from $\mathcal{P}_{a}$, then they will envy any other agent $b$ who gets allocated a house from $\mathcal{P}_{a}$. The objective is the find an allocation of houses to agents that minimizes the number of envious agents.
Ideally, we would like to obtain an allocation of houses to agents such that no agent is envious. Unfortunately, envy-free allocations are not guaranteed to exist, even when the number of agents and the number of houses are equal.
Notice that the House Allocation problem assumes that agents have global information about all other agents, which may be impractical in the real world. It is more likely that agents have information about the allocation of other agents that they know of, like their friends or co-workers. Thus, in this paper we study a generalization of the House Allocation problem, called the Optimal House Allocation of Agent Network (also referred to as the Graphical House Allocation problem in literature). In this problem, we have an agent graph $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$. In our variant, an agent $a$ is only envious of an agent $b$ if all three of the following conditions are satisfied: (i) $a$ is not allocated a house from $\mathcal{P}_{a}$, (ii) $b$ is allocated a house from $\mathcal{P}_{a}$ and (iii) $b$ is a neighbour of $a$ in the graph $\mathcal{G}_{\mathcal{A}}$. Now, the question is to find an allocation that minimizes the number of envious agents. Note that when the agent graph $\mathcal{G}_{\mathcal{A}}$ is a complete graph then Optimal House Allocation of Agent Network is exactly the House Allocation problem.
There is also a notion of happy agents in allocation problems. An agent $a$ is happy when they are allocated a house from $\mathcal{P}_{a}$. Note that an agent $a$ may both be non-envious of any other agent as well as unhappy. This could happen if no house in $\mathcal{P}_{a}$ is allocated to any agent in the closed neighbourhood of $a$ in the graph $\mathcal{G}_{\mathcal{A}}$. While reduction of envy is important, it is also important to maximize the number of agents that are happy. A concrete application arises in university dormitory assignments, where students submit room preferences and are socially connected through shared courses or clubs. In such settings, it is crucial to limit perceived unfairness among friends while ensuring that as many students as possible are assigned to desirable rooms. This dual criterion captures realistic concerns of both equity and satisfaction in constrained allocation environments. Therefore, our second problem, Optimally Happy House Allocation of Agent Network asks for an allocation that, among all minimum envy allocations, has the maximum number of happy agents.
It is easy to see that any NP-hardness result of Optimal House Allocation of Agent Network also translates to NP-hardness of Optimally Happy House Allocation of Agent Network. Also, any algorithmic result of Optimally Happy House Allocation of Agent Network translates to an algorithm for Optimal House Allocation of Agent Network. Now, we describe our results.
1.1 Contributions
First, we consider both Optimal House Allocation of Agent Network and Optimally Happy House Allocation of Agent Network when each agent prefers exactly one house. In other words, for each agent $a$, $|\mathcal{P}_{a}|=1$. We design polynomial time algorithms for both problems under this variant. The details of this are discussed in Section 3, as Theorem 3.1.
Next, we show that even when the preference lists of each agent is at most $2$, the problem becomes $\mathsf{NP}$-hard. It is to be noted that Madathil et. al [15] proved NP-hardness for House Allocation (Optimal House Allocation of Agent Network on agent graphs which are complete graphs) when the preference list of each agent is at most $2$. Our NP-hardness result for Optimal House Allocation of Agent Network holds even when the agent graph is a $3$-regular graph. The same NP-hardness result shows that Optimal House Allocation of Agent Network, and consequently Optimally Happy House Allocation of Agent Network, is $\mathsf{NP}$-hard even when the number $|\mathcal{A}|=n$ of agents is equal to the number $|\mathcal{H}|=m$ of houses. This is in stark contrast to the House Allocation problem, which can be solved in polynomial time when $n=m$. These proofs appear in detail as Theorem 3.5 and Theorem 3.7, in Section 3.
Thus, we turn to exact algorithms for Optimal House Allocation of Agent Network and Optimally Happy House Allocation of Agent Network. Note that there is a trivial brute force algorithm that runs in time $2^{\mathcal{O}(n\log m)}$, where $n$ is the number of agents and $m$ is the number of houses. First, we design an algorithm for Optimal House Allocation of Agent Network that on input $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$ and $\mathcal{H}$ solves the problem in time $2^{|\mathcal{A}|+2|E_{\mathcal{A}}|}\cdot(|\mathcal{A}|+|\mathcal{H}|)^{\mathcal{O}(1)}$. This algorithm performs better than the naive algorithm for all graphs when $|E_{\mathcal{A}}|=o(|\mathcal{A}|\log|\mathcal{H}|)$. The details of this algorithm can be found in Theorem 4.1, Section 4.
Next, we look at other structural properties of the agent graph and design exact algorithms that utilize the structure. We design an alternate algorithm that has the best running time when the agent graph comes from a class of graphs that have small balanced separators. For example, this algorithm is the most efficient when the agent graph is a planar graph or a bounded genus graph. The details of this algorithm can be found in Theorem 4.6, Section 4.
We further design an exact exponential algorithm for Optimal House Allocation of Agent Network for agent graphs with bounded vertex cover size. If the size of the vertex cover of the agent graph is $(\log n)^{\mathcal{O}(1)}$ then this algorithm is the fastest algorithm. On the other hand, we also show that Optimal House Allocation of Agent Network is $\mathsf{NP}$-hard even when the agent graph is a split graph or a complete bipartite graph with maximum vertex cover size at most $n^{\varepsilon}$, where $n$ is the number of agents and $\varepsilon\in(0,1)$ is a constant - therefore, we do not expect to obtain quasi-polynomial time algorithms for the problem when the vertex cover size of the agent graph is $n^{\varepsilon}$.
1.2 Related Work
The House Allocation problem and its various extensions have been extensively studied across economics and computer science [2, 3, 5, 7, 10, 12], particularly in contexts that emphasize fairness and efficiency. We discuss the most relevant prior works to our setting.
Abdulkadiroğlu and Sönmez [1] introduce the House Allocation with Existing Tenants model, which captures scenarios where some agents already occupy houses and may choose to keep them or participate in reallocation. They propose the Top Trading Cycles (TTC) mechanism, which guarantees allocations that are Pareto efficient, individually rational, and strategy-proof. Their model generalizes the classical house allocation setting by incorporating endowments and offering mechanisms that incentivize participation without loss. Focusing on the classical House Allocation problem under ordinal preferences, Gan, Suksompong, and Voudouris [6] develop a polynomial-time algorithm to determine whether an envy-free allocation exists when each agent ranks houses. They explore conditions under which envy-free allocations can or cannot be guaranteed and highlight cases where approximations or partial allocations become necessary. Their work contributes foundational results on the computational aspects of envy-freeness under structured preference inputs and motivates the need for relaxed fairness notions studied in later works.
Hosseini et al. [11, 9] introduce the Graphical House Allocation model, wherein agents are vertices in a graph, and envy is permitted only along the graph’s edges. Assuming identical valuation functions across agents, they focus on minimizing the total envy, measured as the sum of envy across edges. They show that the problem generalizes the classical linear arrangement problem and establish hardness results even for simple graphs like disjoint paths and stars. To address tractability, they develop fixed-parameter algorithms based on graph properties like separability and provide polynomial-time solutions for special families such as paths and cycles. Choo, Ling, Suksompong, Teh, and Zhang [4] address the challenge of achieving envy-freeness in settings where such allocations may not naturally exist. They propose the use of subsidies—monetary compensations that accompany house allocations—to compute envy-free outcomes while minimizing the total subsidy required. They show that the problem is NP-hard in general but solvable in polynomial time when the number of houses differs from the number of agents by a constant, or when agents have identical utilities.
Madathil, Misra, and Sethia. [15] provide a comprehensive study on minimizing envy in House Allocation, introducing three formal measures of envy—number of envious agents, maximum envy, and total envy. They define and analyze the corresponding computational problems (Optimal House Allocation, Egalitarian House Allocation, and Utilitarian House Allocation), establish strong hardness results even under restricted settings like binary preferences and bounded degrees, and design polynomial-time algorithms for structured instances such as extremal preferences. They also present fixed-parameter tractable algorithms based on agent and house types, formulate ILP models with empirical evaluations, and study the trade-off between fairness and welfare through the Price of Fairness, offering tight bounds and identifying cases where fairness incurs no welfare loss.
2 Preliminaries
Let $\mathcal{A}$ be a set of $n$ agents and $\mathcal{H}$ be a set of $m$ houses. Each agent $a\in\mathcal{A}$ is associated with a set of preferred houses $\mathcal{P}_{a}\subseteq\mathcal{H}$. Agent $a$ prefers houses in $\mathcal{P}_{a}$ while disliking the rest. A house allocation is a function $\phi:\mathcal{A}\rightarrow\mathcal{H}$ that is injective, meaning no two agents are assigned the same house. In this paper, $|\mathcal{A}|$ and $|\mathcal{H}|$ shall also be denoted by $n$ and $m$, respectively.
We denote by $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$ a graph over the agent set $\mathcal{A}$, modeling the social or informational structure among agents. An agent $\mathcal{A}$ is said to envy a neighbour $a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)$ under an allocation $\phi$ if $\phi(a)\notin\mathcal{P}_{a}$ and $\phi(a^{\prime})\in\mathcal{P}_{a}$; that is, $\mathcal{A}$ does not receive a house they like, while their neighbour does. Let $\mathcal{E}_{\mathcal{A}}^{\phi}(a)$ denote the set of agents in $N_{\mathcal{G}_{\mathcal{A}}}(a)$ whom agent $\mathcal{A}$ envies under allocation $\phi$, i.e.,
$\mathcal{E}_{\mathcal{A}}^{\phi}(a)=\{a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)\mid\phi(a)\notin\mathcal{P}_{a}\text{ and }\phi(a^{\prime})\in\mathcal{P}_{a}\}.$
We call an agent $a\in\mathcal{A}$ envious if $\left|\mathcal{E}_{\mathcal{A}}^{\phi}(a)\right|\geqslant 1$, otherwise, $a$ is said to be non-envious or envy-free. We call an agent $a\in\mathcal{A}$ happy, if $\phi(a)\in\mathcal{P}_{a}$. The envy (resp. happiness) of the allocation $\phi$ is the total number of envious (resp. happy) agents for the allocation $\phi$. A house is called a dummy house, if it is not preferred by any agent; otherwise we call it a non-dummy house.
We will use the notation $[t]$ to denote the set $\{1,2,\ldots,t\}$. For any function $f:X\to Y$, for a subset $X^{\prime}\subseteq X$, we denote by $f(X^{\prime})$ the set of images of $X^{\prime}$, i.e. $f(X^{\prime})=\{f(x)\mid x\in X^{\prime}\}$.
We are interested in computing allocations $\phi$ that minimize various measures of envy in this graphical setting. Below, we define two optimization problems under this natural perspective capturing envy across agent relationships (defined by $\mathcal{G}_{\mathcal{A}}$). For completeness, we define each of these problems as follows:
Optimal House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes |$\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$.
Optimally Happy House Allocation of Agent Network Input: A set $\mathcal{A}$ of $n$ agents, a graph $\mathcal{G}_{\mathcal{A}}$ on $\mathcal{A}$, a set $\mathcal{H}$ of $m$ houses, and agent preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$. Question: Compute an allocation $\phi$ that minimizes |$\left|\{a\in\mathcal{A}\mid\mathcal{E}_{\mathcal{A}}^{\phi}(a)\neq\emptyset\}\right|$, and among all such minimum-envy allocations, maximizes |$\left|\{a\in\mathcal{A}\mid\phi(a)\in\mathcal{P}_{a}\}\right|$.
Note that Optimally Happy House Allocation of Agent Network is at least as hard as Optimal House Allocation of Agent Network, as any solution to Optimally Happy House Allocation of Agent Network is also a solution to Optimal House Allocation of Agent Network. Whenever we provide algorithms, we analyze this for both Optimal House Allocation of Agent Network and Optimally Happy House Allocation of Agent Network. However, for hardness results, we will only focus on Optimal House Allocation of Agent Network, as these results would automatically translate to Optimally Happy House Allocation of Agent Network as well.
Unless otherwise stated, we assume that preferences are represented extensionally by the preferred set $\mathcal{P}_{a}$ for each agent $\mathcal{A}$, and that all agents receive exactly one, distinct, house. We also assume that envy is only locally defined over $\mathcal{G}_{\mathcal{A}}$, capturing realistic scenarios where agents can compare only within their social or spatial neighbourhoods. Throughout the paper, we use the following additional notation. For any agent $a\in\mathcal{A}$, we write $\deg(a)$ to denote the degree of $\mathcal{A}$ in the graph $\mathcal{G}_{\mathcal{A}}$. We denote by $d=\max\limits_{a\in\mathcal{A}}|\mathcal{P}_{a}|$, the maximum number of houses that a single agent prefers. An allocation is said to be zero-envy if no agent envies any of their neighbours, i.e., $\mathcal{E}_{A}^{\phi}(a)=\emptyset$ for all $a\in\mathcal{A}$. Finally, for any house $h\in\mathcal{H}$, we use $\phi^{-1}(h)$ to denote the agent assigned to house $\mathcal{H}$ under $\phi$, if any.
3 Introducing an Agent Graph: Contrast with the Classical Setting
Notice that the House Allocation problem studied by Madathil et. al [15] is a special case of Optimal House Allocation of Agent Network where the agent graph is a complete graph. In this section, we provide results for Optimal House Allocation of Agent Network that depict that this has a much more interesting complexity landscape in comparison to the House Allocation problem.
We first look into a tractable variant of House Allocation that remains tractable for Optimal House Allocation of Agent Network. Madathil et. al [15] show that when each agent prefers exactly one house, Optimal House Allocation of Agent Network on complete graphs is polynomial-time solvable. We start by extending this result to solving Optimal House Allocation of Agent Network on arbitrary graphs in polynomial time, when each agent prefers exactly one house i.e. $d=1$.
Theorem 3.1.
There is a polynomial-time algorithm for Optimal House Allocation of Agent Network, when every agent prefers exactly one house ($d=1$).
Proof 3.2.
Consider the instance $(\mathcal{A},\mathcal{H},\mathcal{G}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ of Optimal House Allocation of Agent Network. Since each agent $a\in\mathcal{A}$ likes exactly one house, $a$ can be envious of at most one other agent. That is, $|\mathcal{E}_{\mathcal{A}}^{\phi}(a)|\leqslant 1$. We will crucially use this observation to construct a polynomial-time algorithm.
Allocating house $h\in\mathcal{H}$ to agent $a\in\mathcal{A}$ causes the set of agents $Y_{a,h}=\{a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)\mid\mathcal{P}_{a^{\prime}}=\{h\}\}$ to be envious of $a$. Note that in such a case, the agents in $Y_{a,h}$ are not envious of any other agents. This observation would be crucially used to avoid double counting of envious agents.
We now construct a weighted complete bipartite graph $\mathcal{G}_{\mathcal{F}}$ with vertex partitions $\mathcal{A}$ and $\mathcal{H}$. The weights $w(a,h)$ for $a\in\mathcal{A}$, $h\in\mathcal{H}$ are defined as $w(a,h)=|Y_{a,h}|$. The motivation behind defining such a weight function is to model through $w(a,h)$, the number of agents that become envious on assigning house $h\in\mathcal{H}$ to agent $a\in\mathcal{A}$. We finally use a polynomial-time minimum cost maximum bipartite matching algorithm to get a minimum-envy assignment. We formally prove this as follows.
Claim 1.
A minimum-envy allocation has the number of envious agents equal to the cost of the minimum cost maximum matching of $\mathcal{G}_{\mathcal{F}}$. The matching itself corresponds to such an allocation.
Proof 3.3 (Proof of Claim 1).
It suffices to show that every maximum matching in $\mathcal{G}_{\mathcal{F}}$ has a bijective relation with a allocation of houses to agents, where the cost of the matching is equal to the number of envious agents in the allocation.
Let $M$ be a maximum matching; since $|\mathcal{A}|\leqslant|\mathcal{H}|$, each agent must be matched to exactly one house. Define the allocation $\phi$ which assigns each agent $a$, its matched house in $\mathcal{H}$. The weight of any matched edge $\{a,h\}$ denotes the number of envious agents due to assigning $h$ to $a$. Moreover, since each agent can be envious of at most one other agent, the total number of envious agents is simply the sum of the weights of the matching edges: the cost of the matching $M$.
On the other hand, if $\phi$ is an allocation, then consider a matching $M=\{\{a,\phi(a)\}\mid a\in\mathcal{A}\}$ defined by this allocation. This is a matching as no two agents are allocated the same house. If an agent $a^{\prime}$ is envious in $\phi$, it is envious of exactly one of its neighbours, $a\in N_{\mathcal{G}_{\mathcal{A}}}(a^{\prime})$. This happens when $\mathcal{P}_{a^{\prime}}=\{h\}$ and $\phi(a)=h$. Therefore, $a^{\prime}\in Y_{a,h}$ and hence this contributes exactly $1$ to the weight $w(a,\phi(a))$ while not affecting any other matching edge. Indeed, every such envious agent contributes exactly $1$ to exactly one of the matching edges. This again implies that the total number of envious agents in $\phi$ is same as the cost of the matching $M$.
This completes the proof of Claim 1.
Thus we are done.
For completeness, we provide the pseudocode of this polynomial-time algorithm, called Algorithm 1.
Algorithm 1 Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}}$ Output: Minimum envy
1:for $a\in\mathcal{A}$, $h\in\mathcal{H}$ do
2:     $w(a,h)\leftarrow|\{a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)\mid\mathcal{P}_{a^{\prime}}=\{h\}\}|$
3:end for
4:$\mathcal{G}_{\mathcal{F}}\leftarrow$ complete bipartite graph with partitions $\mathcal{A}$, $\mathcal{H}$, and weights $w$
5:return Minimum cost maximum matching of $\mathcal{G}_{\mathcal{F}}$
Note that Algorithm 1 can be modified to solve Optimally Happy House Allocation of Agent Network just by altering the definition of $w(a,h)$ to:
$w(a,h)=\begin{cases}|\{a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)\mid\mathcal{P}_{a^{\prime}}=\{h\}\}|-\frac{1}{n+1}&\text{if }h\in\mathcal{P}_{a}\\ |\{a^{\prime}\in N_{\mathcal{G}_{\mathcal{A}}}(a)\mid\mathcal{P}_{a^{\prime}}=\{h\}\}|&\text{if }h\notin\mathcal{P}_{a}\end{cases}$
A matching in such a bipartite graph will have a cost of $\left(\alpha-\beta\cdot\frac{1}{n+1}\right)$, where $\alpha,\beta\in[n]$ if it corresponds to an allocation with envy $\alpha$ and happiness $\beta$.
Corollary 3.4.
There is a polynomial-time algorithm for Optimally Happy House Allocation of Agent Network, when every agent likes exactly one house ($d=1$).
Now we study Optimal House Allocation of Agent Network when the input graph is restricted to certain simple graph classes. For instance, in the case of independent sets, Optimal House Allocation of Agent Network is trivially solvable as all allocations are zero-envy, and Optimally Happy House Allocation of Agent Network is simply equivalent to the maximum cardinality bipartite matching problem. Moreover, restricting to some arbitrary graph class might potentially increase or decrease the hardness of the problem relative to when the input graph is a complete graph. Hence it is interesting to explore the complexity of Optimal House Allocation of Agent Network when the input graph is restricted to certain simpler graph classes.
Madathil et. al. [15] show that Optimal House Allocation of Agent Network for cliques is hard even when each agent prefers at most two houses, i.e. $d=2$. We show that even for the class of bipartite graphs, Optimal House Allocation of Agent Network remains hard even when each agent prefers at most two houses, i.e. $d=2$. This involves a reduction from the $\mathsf{NP}$-hard problem of CLIQUE, which is a nontrivial extension of the work by Madathil et. al. [15].
Theorem 3.5.
Given an agent graph $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$ and a set $\mathcal{H}$ of houses, Optimal House Allocation of Agent Network is $\mathsf{NP}$-hard, even when $\mathcal{G}_{\mathcal{A}}$ is a complete bipartite graph where each agent prefers at most two houses.
(a) Instance of CLIQUE for $3$-regular graphs, $k=3$
(b) The reduced instance as per Theorem 3.5
Figure 1: Reduction for Theorem 3.5
Proof 3.6.
We provide a reduction from the $\mathsf{NP}$-hard problem of CLIQUE on regular graphs to Optimal House Allocation of Agent Network. Let $(G(V,E),k)$ be an instance of CLIQUE on regular graphs where $G$ is $\delta$-regular and the problem asks to decide if $G$ has a clique of size $k$. Let $|V|=N$ and $|E|=M$.
We reduce $(G(V,E),k)$ to an instance of Optimal House Allocation of Agent Network as follows.
[…]
    Let $\mathcal{A}=\{a^{j}_{v}\mid v\in V,j\in[\delta]\}\cup\{a_{e}\mid e\in E\}$. That is, we have an agent $a_{e}$ for every edge $e$ of $G$, and $\delta$ agents $a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}$ for each vertex $v$ of $G$. Therefore, $n=|\mathcal{A}|=\delta N+M$.
[…]
    Let =$\mathcal{H}=\{h_{v}\mid v\in V\}\cup\{h^{*}_{j}\mid j\in[\delta N+M-k]\}$. Therefore, $|\mathcal{H}|=N+\delta N+M-k=|\mathcal{A}|+(N-k)$.
[…]
    We are now going to set the houses $h_{v}$ to be preferred by agents $a^{j}_{v}$ for all $j\in[\delta]$, and also by $a_{e}$, for all edges $e$ which are incident on $v$. This is effectively done by setting $\mathcal{P}_{a^{j}_{v}}=\{h_{v}\}$, and $\mathcal{P}_{a_{e}}=\{h_{u},h_{v}\}$, where $e=\{u,v\}$. The houses $\{h^{*}_{j}\mid j\in[\delta N+M-k]\}$ are preferred by no agents, i.e., these are dummy houses.
[…]
    We now define the underlying graph $\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}})$. Let $\mathcal{A}_{V}=\{a^{j}_{v}\mid v\in V,j\in[\delta]\}$ and let $\mathcal{A}_{E}=\{a_{e}\mid e\in E\}$. Then $\mathcal{G}_{\mathcal{A}}$ defines the complete bipartite graph between $\mathcal{A}_{V}$ and $\mathcal{A}_{E}$. That is, $E_{\mathcal{A}}=\{\{a,a^{\prime}\}\mid a\in\mathcal{A}_{V},a^{\prime}\in\mathcal{A}_{E}\}$.
[…]
    The target number of envious agents is $k\delta-\binom{k}{2}$. Not that for non-trivial instances, $k\leqslant\delta$, hence $k\delta-\binom{k}{2}$ is a positive integer.
Please refer to Figure 1 for clarity. Every step can be done in polynomial time; hence this is a polynomial-time reduction. We now show the correctness of the reduction.
[…]
    Assign $h_{v}$ to $a^{1}_{v}$ for all $v\in S$.
[…]
    For the remaining $|\mathcal{A}|-k=\delta N+M-k$ agents, assign them a dummy house each.
In $\phi$ the houses $h_{v}$ are left unassigned for $v\notin S$. No agent in $\mathcal{A}_{V}$ are envious as all their neighbours (i.e. $\mathcal{A}_{E}$) are assigned dummy houses. Moreover, $a_{e}$ is not envious if $e$ is not incident on $S$; this is because the houses preferred by $a_{e}$ are unallocated. Therefore, the number of envious agents is at most the number of edges incident on $S$; this is precisely $k\delta-\binom{k}{2}$, as $G$ is $\delta$-regular and $S$ is a clique of size $k$.
Reverse direction. Let $\phi$ be an allocation with at most $k\delta-\binom{k}{2}$ envious agents. We say that an allocation $\phi$ is nice if for all $v\in V$, the house $h_{v}$ is either unassigned, or assigned to the agent $a^{1}_{v}$. If $\phi$ is not nice to begin with, then $\phi$ can be converted to a nice allocation by a sequence of exchanges, none of which increase total envy.
Let $h_{v}$ be a house that is assigned to an agent in $\mathcal{A}$ other than $a^{1}_{v}$.
[…]
    Case I: $h_{v}$ is assigned to $a^{j}_{v}$ for some $j\neq 1$. We swap the houses allocated to $a^{j}_{v}$ and $a^{1}_{v}$. This does not change the number of envious agents, due to the symmetry of $\mathcal{G}_{\mathcal{A}}$.
[…]
    Case II: $h_{v}$ is assigned to $a^{j}_{u}$ for some $j\in[\delta]$, $u\neq v$. We swap the houses allocated to $a^{j}_{u}$ and $a^{1}_{v}$. Again, this does not increase the number of envious agents. This is because
    1. 1.
      $a^{1}_{v}$ is not envious in both allocations.
    2. 2.
      $a^{j}_{u}$ did not have their preferred house before the swap. Since the neighbourhood of $a^{j}_{u}$ is unaffected by the swap, $a^{j}_{u}$ cannot become envious after the swap if $a^{j}_{u}$ was not envious before the swap.
    3. 3.
      For all other agents, they are envious after the swap if and only if they were envious before the swap. This is because the set of houses allocated to their neighbours remain unchanged.
[…]
    Case III: $h_{v}$ is assigned to $a_{e}$ for some $e\in E$. $a^{1}_{v},a^{2}_{v}\ldots,a^{\delta}_{v}$ must be envious of $a_{e}$ in $\phi$. We now look at the sequence of agents $a^{1}_{v_{0}}=a^{1}_{v},a^{1}_{v_{1}},\ldots,a^{1}_{v_{p-1}},a^{1}_{v_{p}}$ such that
    • $\phi(a_{e})=h_{v_{0}}$.
    • $\phi(a^{1}_{v_{0}})=h_{v_{1}},\phi(a^{1}_{v_{1}})=h_{v_{2}},\ldots,\phi(a^{1}_{v_{p-1}})=h_{v_{p}}$.
    • $\phi(a^{1}_{v_{p}})=h^{*}_{q}$ is a dummy house.
    Note that such a sequence must exist and end with an agent who is assigned a dummy house, as the number of agents and houses are finite. Therefore, $p$ is a finite integer. Now, we modify the allocation to $\phi^{\prime}$ as follows:
    • $\phi^{\prime}(a_{e})=h^{*}_{q}$.
    • $\phi^{\prime}(a^{1}_{v_{0}})=h_{v_{0}},\phi^{\prime}(a^{1}_{v_{1}})=h_{v_{1}},\ldots,\phi^{\prime}(a^{1}_{v_{p-1}})=h_{v_{p-1}},\phi^{\prime}(a^{1}_{v_{p}})=h_{v_{p}}$.
    • $\phi^{\prime}(a)=\phi(a)$ for all other agents $a$.
    We will now argue why this set of exchanges do not increase the number of envious agents.
    • Case A: $e$ is incident on $v$. Let $e=e_{1},e_{2},\ldots,e_{\delta}$ be the edges incident on $v$. In $\phi$, $a_{e}$ was not envious, but agents $a_{e_{2}},\ldots,a_{e_{\delta}}$ may or may not be envious. The agents $a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}$ were all envious in $\phi$. In $\phi^{\prime}$, $a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}$ are not envious, agent $a_{e}$ becomes envious, and agents $a_{e_{2}},\ldots,a_{e_{\delta}}$ may become envious. Therefore, the number of envious agents in $\{a_{e},a_{e_{2}},a_{e_{3}},\ldots,a_{e_{\delta}}\}\cup\{a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}\}$ do not increase. The agents $a^{1}_{v_{1}},a^{1}_{v_{2}},\ldots,a^{1}_{v_{p}}$ all become non-envious in $\phi^{\prime}$ as they get their preferred houses. For all other agents $a^{j}_{u}$ where $u$ is a vertex in $V$, the dummy house $h^{*}_{q}$ is the only house which appears in the set of houses assigned to neighbours of $a^{j}_{u}$ in $\phi^{\prime}$ which did not appear in $\phi$; hence $a^{j}_{u}$ is envious in $\phi^{\prime}$ only if they were envious in $\phi$. Finally, all other agents $a_{e^{\prime}}$ where $e^{\prime}$ is an edge in $E$, are envious in $\phi^{\prime}$ if and only if they were envious in $\phi$ as only $h_{v}$ (a house not preferred by $a_{e^{\prime}}$) gets added to the set of houses assigned to their neighbours.
    • Case B: $e$ is not incident on $v$. Let $e_{1},e_{2},\ldots,e_{\delta}$ be the edges incident on $v$. In $\phi$, the agents $a_{e_{1}},a_{e_{2}},\ldots,a_{e_{\delta}}$ may or may not be envious; but the agents $a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}$ were all envious in $\phi$. In $\phi^{\prime}$, $a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}$ are not envious, while agents $a_{e_{1}},a_{e_{2}},\ldots,a_{e_{\delta}}$ may become envious. Therefore, the number of envious agents in $\{a_{e_{1}},a_{e_{2}},a_{e_{3}},\ldots,a_{e_{\delta}}\}\cup\{a^{1}_{v},a^{2}_{v},\ldots,a^{\delta}_{v}\}$ do not increase. The agents $a^{1}_{v_{1}},a^{1}_{v_{2}},\ldots,a^{1}_{v_{p}}$ all become non-envious in $\phi^{\prime}$ as they get their preferred houses. If the agent $a_{e}$ is not envious in $\phi$, then it cannot become envious in $\phi^{\prime}$ as only $h_{v}$ (a house not preferred by $a_{e}$) gets added to the set of houses assigned to its neighbours. For all other agents $a^{j}_{u}$ where $u$ is a vertex in $V$, the dummy house $h^{*}_{q}$ is the only house which appears in the set of houses assigned to neighbours of $a^{j}_{u}$ in $\phi^{\prime}$ which did not appear in $\phi$; hence $a^{j}_{u}$ is envious in $\phi^{\prime}$ only if they were envious in $\phi$. Finally, all other agents $a_{e^{\prime}}$ where $e^{\prime}$ is an edge in $E$, are envious in $\phi^{\prime}$ if and only if they were envious in $\phi$ as only $h_{v}$ (a house not preferred by $a_{e^{\prime}}$) gets added to the set of houses assigned to their neighbours.
These exchanges do not increase the total envy, but they allocate at least one house $h_{v}$ to $a^{1}_{v}$, which was previously allocated to some other agent. Repeating this for at most $|\mathcal{H}|$ times would give us a nice allocation, with envy no more than that in $\phi$.
Hence, we can safely assume that $\phi$ is a nice allocation with envy at most $k\delta-\binom{k}{2}$. $\phi$ allocates dummy houses to $a_{e}$ for all $e\in E$ and to $a^{2}_{v},a^{3}_{v},\ldots,a^{\delta}_{v}$ for all $v\in V$. Moreover, it is safe to assume that all dummy houses are assigned to some agent, otherwise we can assign a dummy house to any $a^{1}_{v}$ without increasing total envy.
We now look into agents which are not assigned dummy vertices by $\phi$. There are $|\mathcal{A}|-(\delta N+M-k)=(\delta N+M)-(\delta N+M-k)=k$ many such agents, all of the form $a^{1}_{v}$, $v\in V$. Let $a^{1}_{v_{1}},a^{1}_{v_{2}},\ldots,a^{1}_{v_{k}}$ be the agents which are not assigned dummy houses. For any $e\in E$, $e$ is incident on least one vertex in $v_{1},v_{2},\ldots,v_{k}$, if and only if the agent $a_{e}$ is envious. Therefore, the number of envious agents is precisely the number of edges which are incident on the vertices $v_{1},v_{2},\ldots,v_{k}$. Let $\gamma$ be the number of edges with both endpoints in $\{v_{1},v_{2},\ldots,v_{k}\}$. The number of edges incident on $\{v_{1},v_{2},\ldots,v_{k}\}$ is therefore equal to $k\delta-\gamma$. By assumption, the number of envious agents is at most $k\delta-\binom{k}{2}$. Thus $\gamma\geqslant\binom{k}{2}$, implying $v_{1},v_{2},\ldots,v_{k}$ is a clique.
This completes the proof of correctness of the reduction.
In Section 5, we further prove that restricting the agent graphs to bipartite graphs (Theorem 5.4) and split graphs (Theorem 5.6) with very small vertex cover sizes still yields $\mathsf{NP}$-hardness.
Moreover, Madathil et. al [15] provide polynomial-time algorithms for the following cases: (i) when $|\mathcal{A}|=|\mathcal{H}|$, (ii) when preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$ follow an extremal structure. Preferences follow an extremal structure if there is an ordering of houses in =$\mathcal{H}=\{h_{1},h_{2},\ldots,h_{m}\}$ such that for every agent $a\in\mathcal{A}$, they prefer either some suffix of the houses, $\mathcal{P}_{a}=\{h_{i},h_{i+1},\ldots,h_{m}\}$, or some prefix of the houses $\mathcal{P}_{a}=\{h_{1},h_{2},\ldots,h_{i}\}$, for some $i\in[m]$.
In contrast, we show that even when we restrict the graph of an Optimal House Allocation of Agent Network instance to be 3-regular, and have $|\mathcal{A}|=|\mathcal{H}|$ and identical preferences for agents, the problem is $\mathsf{NP}$-hard. Note that identical preferences, i.e. $\mathcal{P}_{a}=\mathcal{P}_{a^{\prime}}$ for all $a,a^{\prime}\in\mathcal{A}$, is a special case of extremal structure as we can order the houses where all dummy houses are ordered after all houses preferred by every agent.
Theorem 3.7.
Given an agent graph $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},\mathcal{E}_{\mathcal{A}})$ and a set $\mathcal{H}$ of houses, Optimal House Allocation of Agent Network is $\mathsf{NP}$-hard even when the input graph is a 3-regular graph with $|\mathcal{A}|=|\mathcal{H}|$ and agents have identical preferences (i.e. $\mathcal{P}_{a}=\mathcal{P}_{a^{\prime}}$ for all $a,a^{\prime}\in\mathcal{A}$).
Proof 3.8.
We provide a reduction from the $1/2$-Vertex Separator problem [16] on 3-regular graphs to Optimal House Allocation of Agent Network. The problem takes input a graph $G(V,E)$ and an integer $k$ and asks if there is a subset $S$ of $V$ that $|S|\leqslant k$ such that $V\setminus S$ can be partitioned into two equal-sized subsets $S_{1}$, $S_{2}$ such that no edge connects a vertex in $S_{1}$ with a vertex in $S_{2}$. This problem is shown to be $\mathsf{NP}$-complete even for 3-regular graphs by Müller and Wagner [16].
Let $(G(V,E),k)$ be an instance of $1/2$-Vertex Separator problem such that $G(V,E)$ is 3-regular. Let $V=\{a_{1},a_{2},\ldots,a_{n}\}$. We create an instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},2\left\lfloor\frac{k}{2}\right\rfloor)$ of Optimal House Allocation of Agent Network as follows:
[…]
    $\mathcal{A}=V=\{a_{1},a_{2},\ldots,a_{n}\}$.
[…]
    =$\mathcal{H}=\{h_{1},h_{2},\ldots,h_{n}\}$.
[…]
    $E_{\mathcal{A}}=E$.
[…]
    Let =$\mathcal{H}^{*}=\{h_{1},h_{2},\ldots,h_{t}\}$, where $t=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$. $\mathcal{P}_{a}=\mathcal{H}^{*}$ for all $a\in\mathcal{A}$.
Intuitively the underlying agent graph is the same as $G$, and every agent prefers the same $t=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$ houses. Note that $t$ is an integer since the number of vertices of $G$, $n$, is even as $G$ is 3-regular. This reduction can be computed in polynomial time. We now show the correctness of the reduction.
Claim 2.
$(G,k)$ and $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ are equivalent instances of the $1/2$-Vertex Separator problem, if $G$ is 3-regular.
Proof 3.9.
This holds trivially when $k$ is even. Further, if $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance, then so is $(G,k)$ as $k\geqslant 2\left\lfloor\frac{k}{2}\right\rfloor$.
Now, say $k$ is odd. and let $(G,k)$ be a Yes-instance. Let $k=2\lambda+1$ for $\lambda\in\mathbb{Z}$. Note that since $G$ is 3-regular, $n=|V|$ must be even. Let $S$ be any separator that separates $V\setminus S$ into equal sized, disjoint, and non-adjacent parts $S_{1}$, $S_{2}$ and satisfying $|S|\leqslant k$. Therefore, $|S_{1}|=|S_{2}|$. Hence $|V|-|S|=|S_{1}|+|S_{2}|$, or $|S|=n-2|S_{1}|$, an even integer: say $|S|=2\beta$, for some integer $\beta$. Therefore, $0.52\beta\leqslant 2\lambda+1\implies\beta\leqslant\lambda+$ 0.5. Since $\beta,\lambda$ are integers, we must have $\beta\leqslant\lambda$; implying $|S|=2\beta\leqslant 2\lambda=2\left\lfloor\frac{k}{2}\right\rfloor$. Hence $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ is also a Yes-instance.
Claim 3.
If $(G(V,E),2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance of the $1/2$-Vertex Separator problem, where $G$ is 3-regular, there exists $S\subseteq V$ of size $|S|=2\left\lfloor\frac{k}{2}\right\rfloor$ that separates $V\setminus S$ into equal sized, disjoint, and non-adjacent parts.
Proof 3.10.
Let $S^{\prime}$ be any separator that separates $V\setminus S$ into two equal sized, disjoint, and non-adjacent parts $S_{1}^{\prime},S_{2}^{\prime}$, such that $|S^{\prime}|\leqslant 2\left\lfloor\frac{k}{2}\right\rfloor$. Such an $S^{\prime}$ exists as $(G(V,E),2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance.
Let $S_{1}^{\prime}=\{x_{1},x_{2},\ldots,x_{s}\}$ and $S_{2}^{\prime}=\{y_{1},y_{2},\ldots,y_{s}\}$ for some $s$, where $x_{i}$ is not adjacent to $y_{j}$ for all $i,j\in[s]$. Therefore, $|S^{\prime}|=n-2s$, an even number. This gives us
$n-2s\leqslant 2\left\lfloor\frac{k}{2}\right\rfloor\implies s-\frac{n}{2}+\left\lfloor\frac{k}{2}\right\rfloor\geqslant 0$
Define $\gamma=s-\frac{n}{2}+\left\lfloor\frac{k}{2}\right\rfloor$, note that $0\leqslant\gamma\leqslant s$. We now construct our target separator $S$.
[…]
    define $S_{1}=\{x_{\gamma+1},x_{\gamma+2},\ldots,x_{s}\}$
[…]
    define $S_{2}=\{y_{\gamma+1},y_{\gamma+2},\ldots,y_{s}\}$
[…]
    define $S=S^{\prime}\cup\{x_{1},x_{2},\ldots,x_{\gamma}\}\cup\{y_{1},y_{2},\ldots,y_{\gamma}\}$
We have $S\cup S_{1}\cup S_{2}=V$, and $S$ separators $V\setminus S$ into equal sized, disjoint, and non-adjacent subsets $S_{1}$ and $S_{2}$. Moreover,
$|S|=|S^{\prime}|+2\gamma=(n-2s)+2\left(s-\frac{n}{2}+\left\lfloor\frac{k}{2}\right\rfloor\right)=2\left\lfloor\frac{k}{2}\right\rfloor$
This completes the proof of Claim 3.
[…]
Observe that no edge in $E_{\mathcal{A}}$ can connect an agent in $X$ to an agent in $Y$; otherwise the corresponding agent in $Y$ would be envious of the corresponding agent in $X$. Note that $|Y\cup J|=\frac{n}{2}+\left\lfloor\frac{k}{2}\right\rfloor\geqslant 2\left\lfloor\frac{k}{2}\right\rfloor$. Let $J^{\prime}$ be an arbitrary subset of $Y\cup J$ and a superset of $J$ satisfying $|J^{\prime}|=2\left\lfloor\frac{k}{2}\right\rfloor$. Then $J^{\prime}$ is separates $V=\mathcal{A}$ into $X$ and $(Y\setminus J^{\prime})$ satisfying $|X|=|Y\setminus J^{\prime}|=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$. Therefore, $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance of the $1/2$-Vertex Separator problem, and so is $(G,k)$ (by Claim 2).
Forward direction. On the other hand, if $(G,k)$ is a Yes-instance for the $1/2$-Vertex Separator problem, then so is $(G,2\left\lfloor\frac{k}{2}\right\rfloor)$ (by Claim 2). By Claim 3, there exists a partition $S,X,Y$ of $\mathcal{A}$ such that $|S|=2\left\lfloor\frac{k}{2}\right\rfloor$, $|X|=|Y|=\frac{n}{2}-\left\lfloor\frac{k}{2}\right\rfloor$, and there is no edge between $X$ and $Y$ in $\mathcal{G}_{\mathcal{A}}$. We create an allotment $\phi$ defined as follows:
[…]
    Assign all houses in $\mathcal{H}^{*}$ to agents in $X$ (arbitrarily).
[…]
    Assign all other houses to all other agents (arbitrarily).
The only envious agents can be the ones in $S$. Hence there are at most $2\left\lfloor\frac{k}{2}\right\rfloor$ many envious agents. Therefore, the Optimal House Allocation of Agent Network instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},2\left\lfloor\frac{k}{2}\right\rfloor)$ is a Yes-instance.
This completes the correctness of the reduction.
With this, we derive that even when restricting to certain simple graph classes the problem of Optimal House Allocation of Agent Network might be harder than when restricting to just complete graphs.
4 Exact Algorithms
In this section, we design exact algorithms for Optimal House Allocation of Agent Network. The naive brute force algorithm that enumerates all possible assignments of $m$ houses to $n$ agents takes time $2^{\mathcal{O}(n\log m)}$. We design exact algorithms that are more efficient than this. Each algorithm caters to different graph properties of the input agent graphs and therefore become most efficient in different scenarios depending on the input agent graph. All the algorithms extend easily to work for Optimally Happy House Allocation of Agent Network as well. First, we design an algorithm that is single exponential in the number of edges in the agent graph. This algorithm is more efficient that the above naive algorithm when the agent graph is a sparse graph. We also design an alternate separator-based algorithm that has better running time for graph classes with small balanced separators, like the class of planar graphs and bounded genus-graphs.
4.1 Single Exponential Exact Algorithms
First, we propose an exact algorithm with running time being single exponential in the number of nodes and edges.
Theorem 4.1.
There exists an algorithm that solves Optimal House Allocation of Agent Network for the instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ in time $2^{n+2|E_{\mathcal{A}}|}\cdot(n+m)^{\mathcal{O}(1)}$, where $n=|\mathcal{A}|$ and $m=|\mathcal{H}|$.
Proof 4.2.
Consider the instance $(\mathcal{A},\mathcal{H},\mathcal{G}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$. Agent $a$ is envious on agent $a^{\prime}$ only if $\{a,a^{\prime}\}$ is an edge in $E_{\mathcal{A}}$. The flow of our algorithm will be as follows: (i) guess which agent is envious of which other agent, (ii) among the non-envious agents guess which subset of agent gets their preferred houses, (iii) use a bipartite matching algorithm to decide if such a allotment exists which matches our guesses.
For an allotment $\phi$, recall that $\mathcal{E}_{\phi}^{\mathcal{H}}(a)$ is the set of agents $a^{\prime}$ where $a$ is envious on $a^{\prime}$. If $N_{\mathcal{G}_{\mathcal{A}}}(a)$ is the set of neighbouring agents of $a$, then $\mathcal{E}_{\phi}^{\mathcal{H}}(a)\subseteq N_{\mathcal{G}_{\mathcal{A}}}(a)$. Since $|N_{\mathcal{G}_{\mathcal{A}}}(a)|=\deg(a)$ is the degree of $a$ in $\mathcal{G}_{\mathcal{A}}$, there are $2^{\deg(a)}$ possible choices of $\mathcal{E}_{\phi}^{\mathcal{H}}(a)$. We make a guess of $\mathcal{E}_{\phi}^{\mathcal{H}}(a)$ for all $a\in\mathcal{A}$.
The agents with $\mathcal{E}_{\phi}^{\mathcal{H}}(a)=\emptyset$ are non-envious. A fraction of these agents receive their preferred houses, while all other agents receive a house not preferred by them. We make a guess $C\subseteq\{a\in\mathcal{A}\mid\mathcal{E}_{\phi}^{\mathcal{H}}(a)=\emptyset\}$ of the agents who receive their preferred houses. There are at most $2^{n}$ possible guesses of $C$.
We now fix $C$ and $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$ and we wish to check if there are any allotments which matches such a guess. We maintain a set $\mathcal{F}_{a}\subseteq\mathcal{H}$ for each $a\in\mathcal{A}$ which would denote the houses that can be assigned to $a$ such that it matches the guess of $C$ and $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$. We compute $\mathcal{F}_{a}$ as follows.
[…]
    We start with $\mathcal{F}_{a}=\mathcal{H}$ and iteratively trim $\mathcal{F}_{a}$ by imposing restrictions, one by one.
[…]
    If $a\notin C$, then $\phi(a)\notin\mathcal{P}_{a}$. Thus $\mathcal{F}_{a}$ cannot contain any house in $\mathcal{P}_{a}$. While, for $a\in C$, $\mathcal{F}_{a}$ can only contain houses in $\mathcal{P}_{a}$.
[…]
    For an edge $\{a,a^{\prime}\}$ in $E_{\mathcal{A}}$, if $a\in\mathcal{E}_{\phi}^{\mathcal{H}}(a^{\prime})$, then $\phi(a^{\prime})\notin\mathcal{P}_{a^{\prime}}$ but $\phi(a)\in\mathcal{P}_{a^{\prime}}$. Thus $\mathcal{F}_{a}$ can contain only houses in $\mathcal{P}_{a^{\prime}}$. Note that we do not need to impose any further constraints on $\mathcal{F}_{a^{\prime}}$ as the previous step already trimmed $\mathcal{F}_{a^{\prime}}$ owing to $a\notin C$.
[…]
    Otherwise, for an edge $\{a,a^{\prime}\}$ in $E_{\mathcal{A}}$, if $a\in\mathcal{E}_{\phi}^{\mathcal{H}}(a^{\prime})$ and $a^{\prime}\notin C$, then $\phi(a^{\prime})\notin\mathcal{P}_{a^{\prime}}$ and $\phi(a)\notin\mathcal{P}_{a^{\prime}}$. Therefore, $\mathcal{F}_{a}$ cannot contain any house in $\mathcal{P}_{a^{\prime}}$.
[…]
    Finally, if for $\{a,a^{\prime}\}\in E_{\mathcal{A}}$, if $a\notin\mathcal{E}_{\phi}^{\mathcal{H}}(a^{\prime})$ and $a^{\prime}\in C$, then $\phi(a^{\prime})\in\mathcal{P}_{a^{\prime}}$. Any house can be assigned to $a$ without making $a^{\prime}$ envious; this imposes no further restrictions on $\mathcal{F}_{a}$.
Note that, by construction of $(\mathcal{F}_{a})_{a\in\mathcal{A}}$, the set of allotments $\phi$ that correspond to the guesses $C$, and $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$ are exactly the set of allotments $\phi$, satisfying $\phi(a)\in\mathcal{F}_{a}$. Since all such $\phi$ have the same envy, $\xi=|\{a\in\mathcal{A}\mid\mathcal{E}_{\phi}^{\mathcal{H}}(a)\neq\emptyset\}|$, it suffices to check if at least one $\phi$ exists that correspond to the guesses of $C$ and $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$. This check can be done using any maximum bipartite matching algorithm on the graph defined by $(\mathcal{F}_{a})_{a\in\mathcal{A}}$.
We already argued the correctness of this algorithm, all that remains is to argue its running time. For a fixed guess of $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$ and $C$, the algorithm runs in polynomial time, including the maximum bipartite matching subroutine. There are at most $2^{n}$ guesses of $C$, while there are $2^{\deg(a)}$ guesses of $\mathcal{E}_{\phi}^{\mathcal{H}}(a)$. Therefore, the total number of guesses of $(\mathcal{E}_{\phi}^{\mathcal{H}}(a))_{a\in\mathcal{A}}$ and $C$ are,
$2^{n}\cdot\prod\limits_{a\in A}2^{\deg(a)}=2^{n+2|E_{\mathcal{A}}|}$
Hence the algorithm terminates with a correct output in time $2^{n+2|E_{\mathcal{A}}|}\cdot(n+m)^{\mathcal{O}(1)}$, where $n=|\mathcal{A}|$ and $m=|\mathcal{H}|$.
We name this as Algorithm 2, the pseudocode of which is given below. Note that this algorithm is more efficient than the brute force for graphs with $|E_{\mathcal{A}}|=o(n\log m)$, these include sparse graph classes like planar graphs, bounded treewidth graphs and bounded genus graphs.
Algorithm 2 Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}}$ Output: Minimum envy
1:$\xi\leftarrow n$triangleright In the worst case, all agents are envious
2:for $(\mathcal{E}_{\phi}^{\mathcal{H}})_{a\in\mathcal{A}}\in(2^{N_{\mathcal{G}_{\mathcal{A}}}(a)})_{a\in\mathcal{A}}$ do $\triangleright$ Guess $\mathcal{E}_{\phi}^{\mathcal{H}}(a)\subseteq N_{\mathcal{G}_{\mathcal{A}}}(a)$ for all $a\in\mathcal{A}$
3:     for $C\subseteq\{a\in\mathcal{A}\mid\mathcal{E}_{\phi}^{\mathcal{H}}(a)=\emptyset\}$ do $\triangleright$ Guess $C\subseteq\mathcal{A}$, agents getting preferred house
4:         for $a\in\mathcal{A}$ do
5:              if $a\in C$ then
6:                  $\mathcal{F}_{a}\leftarrow\mathcal{P}_{a}$
7:              else
8:                  $\mathcal{F}_{a}\leftarrow\mathcal{H}\setminus\mathcal{P}_{a}$
9:              end if
10:         end for
11:         for $\{a,a^{\prime}\}\in E_{\mathcal{A}}$ do
12:              if $a\in\mathcal{E}_{\phi}^{\mathcal{H}}(a^{\prime})$ then $\triangleright$ Trim $\mathcal{F}_{a}$
13:                  $\mathcal{F}_{a}\leftarrow\mathcal{F}_{a}\cap\mathcal{P}_{a^{\prime}}$
14:              else if $a^{\prime}\notin C$ then
15:                  $\mathcal{F}_{a}\leftarrow\mathcal{F}_{a}\setminus\mathcal{P}_{a^{\prime}}$
16:              end if
17:              if $a^{\prime}\in\mathcal{E}_{\phi}^{\mathcal{H}}(a)$ then $\triangleright$ Repeat for $a^{\prime}$
18:                  $\mathcal{F}_{a^{\prime}}\leftarrow\mathcal{F}_{a^{\prime}}\cap\mathcal{P}_{a}$
19:              else if $a\notin C$ then
20:                  $\mathcal{F}_{a^{\prime}}\leftarrow\mathcal{F}_{a^{\prime}}\setminus\mathcal{P}_{a}$
21:              end if
22:         end for
23:         $E_{\mathcal{F}}\leftarrow\{\{a,h\}\mid a\in\mathcal{A},h\in\mathcal{F}_{a}$}; $\mathcal{G}_{\mathcal{F}}\leftarrow(\mathcal{H}\cup\mathcal{A},E_{\mathcal{F}})$triangleright Feasibility bipartite graph
24:         if maximum cardinality bipartite matching of $\mathcal{G}_{\mathcal{F}}$ is of size $|\mathcal{A}|$ then
25:              $\xi\leftarrow\min(\xi,|\{a\in\mathcal{A}\mid\mathcal{E}_{\phi}^{\mathcal{H}}(a)\neq\emptyset\}|$)
26:         end if
27:     end for
28:end for
29:return $\xi$
Remark 4.3.
Algorithm 2 can be trivially adapted to output the actual allocation in the same asymptotic runtime. Moreover, the algorithm can also be adapted to solve Optimally Happy House Allocation of Agent Network; this can be done by maximizing $C$.
4.2 Balanced Separator Based Algorithm
The notion of envy is intrinsically linked to the underlying graph structure $\mathcal{G}_{\mathcal{A}}$ and the exact edges play an important role in contributing to the total envy of any allotment. Two agents which do not share an edge in $E_{\mathcal{A}}$ can never be envious of each other. So, it makes sense to look into separators of the graph $\mathcal{G}_{\mathcal{A}}$ which separate $\mathcal{A}$ into nearly equal sized $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ which are independent of each other in the problem.
We start of with a definition of $f(n)$-balanced-separable graph classes.
Definition 4.4.
A graph class $\mathfrak{G}$ is $f(n)$-balanced-separable, if for all graphs $\mathcal{G}\in\mathfrak{G}$ of $n$ vertices,
[…]
    all induced subgraphs of $\mathcal{G}$ are also contained in $\mathfrak{G}$,
[…]
    there exists a partition $(S,V_{1},V_{2})$ of the vertex set of $\mathcal{G}$ such that $|S|\leqslant f(n)$, $|V_{1}|,|V_{2}|\leqslant 2n/3$ and there is no edge in $\mathcal{G}$ with one endpoint in $V_{1}$ and the other in $V_{2}$.
For example, the class of planar graphs are $\mathcal{O}(\sqrt{n})$-balanced-separable (due to the planar separator theorem [14]), graphs of treewidth tw are $(\text{tw}+1)$-balanced-separable.
We look into the problem of Optimal House Allocation of Agent Network when the underlying graph $\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}})$ is from an $f(n)$-balanced-separable graph class. Let $S\subseteq\mathcal{A}$ be a subset of agents of size $f(n)$, such that $\mathcal{A}\setminus S$ can be partitioned into $\mathcal{A}_{1},\mathcal{A}_{2}$ where $|\mathcal{A}_{1}|,|\mathcal{A}_{2}|\leqslant 2n/3$ and there is no edge between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Thus, the exact allotment of houses to agents in $\mathcal{A}_{2}$ would not affect which agents in $\mathcal{A}_{1}$ would be envious. This would allow us to recurse into independent subproblems with $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ as the set of agents.
We define a harder version of Optimal House Allocation of Agent Network, called Annnotated Optimal House Allocation of Agent Network. Firstly, we allow the input to specify a feasibility set $\mathcal{F}_{a}$ for each agent $a\in\mathcal{A}$ which asks to only consider allotments $\phi$ satisfying $\phi(a)\in\mathcal{F}_{a}$. Secondly, there is a provision to mark agents ‘angry’; an angry agent becomes envious whenever they don’t get a house they prefer (irrespective of what their neighbours are assigned). The input specifies a subset $B\subseteq\mathcal{A}$ of agents such that an agent $a\in B$ is envious if $\phi(a)\notin\mathcal{P}_{a}$. As usual, for all other agents $a\in\mathcal{A}\setminus B$, the agent $a$ is envious if $\phi(a)\in\mathcal{P}_{a}$, and there exists $\{a,a^{\prime}\}\in E_{\mathcal{A}}$ such that $\phi(a^{\prime})\in\mathcal{P}_{a}$.
Annnotated Optimal House Allocation of Agent Network Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$. Question: Compute an allocation $\phi$ satisfying $\phi(a)\in\mathcal{F}_{a}$ which minimizes |$|\{a\in\mathcal{A}\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in\mathcal{A},\phi(a^{\prime})\in\mathcal{P}_{a}\}|+|\{a\in B\mid\phi(a)\notin\mathcal{P}_{a}\}|$
Remark 4.5.
Annnotated Optimal House Allocation of Agent Network becomes the problem Optimal House Allocation of Agent Network when $B=\emptyset$, and $\mathcal{F}_{a}=\mathcal{H}$ for all $a\in\mathcal{A}$.
We now propose a recursive algorithm to solve Annnotated Optimal House Allocation of Agent Network using balanced separators of size $f(n)$.
Theorem 4.6.
There exists an algorithm that solves Annnotated Optimal House Allocation of Agent Network, and hence Optimal House Allocation of Agent Network for the instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ in time $2^{m+\mathcal{O}(f(n)\cdot\log^{2}n+n)}$ where $n=|\mathcal{A}|$ and $m=|\mathcal{H}|$ when the input graph is from a $f(n)$-balanced-separable graph class.
Proof 4.7.
Let $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ be an instance of Optimal House Allocation of Agent Network. Firstly, we only look into the case where the number of houses is same as the number of agents, i.e. $n=m$. If that is not the case, we can guess which houses appear in the final allotment, which blows up the final algorithm by a factor of at most $2^{m}$.
Let $S\subseteq\mathcal{A}$ be any subset of agents of size $f(n)$, such that $\mathcal{A}\setminus S$ can be partitioned into $\mathcal{A}_{1},\mathcal{A}_{2}$ where $|\mathcal{A}_{1}|,|\mathcal{A}_{2}|\leqslant 2n/3$ and there is no edge between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. $S$ can be computed naively using a brute force algorithm that enumerates all $\binom{n}{f(n)}$ subsets of size $f(n)$ and checking which of them form balanced separators.
Our next step is to guess the houses that are allocated to agents in $S$, i.e. we guess $\phi(a)\in\mathcal{F}_{a}$ for all $a\in S$. There are at most $n^{f(n)}$ many such guesses. We denote by $\phi(S)$ the set $\{\phi(a)\mid a\in S\}$. For a subset of agents in $S$, $\phi$ might already assign them their preferred houses, call this $C=\{a\in S\mid\phi(a)\in\mathcal{P}_{a}\}$. Some agents in $S$ which are not angry (i.e. not in $B$) might be already envious of some other agent in $S$, call this set $D=\{a\in S\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$. Some other agents who were angry might become envious from not getting a preferred house, call this set $Q=\{a\in S\cap B\mid a\notin\mathcal{P}_{a}\}$. Let the remaining agents in $S$ be $R=S\setminus(C\cup D\cup Q)$.
The agents in $C$ are non-envious, and the agents in $D\cup Q$ are envious regardless of the allocation of the agents in $\mathcal{A}_{1}\cup\mathcal{A}_{2}$. However, the agents in $R$ can be either envious or non-envious depending on what houses are allocated to which agents in $\mathcal{A}_{1}\cup\mathcal{A}_{2}$. Note that $B\cap S\subseteq C\cup Q$, therefore, $R$ does not contain any angry agent in $A\in B$. Agents $a$ in $R$ got assigned some house other than their preference, so they are envious if and only if any of their neighbour in $\mathcal{A}_{1}\cup\mathcal{A}_{2}$ gets assigned a house in $\mathcal{P}_{a}$.
We guess a subset $K\subseteq R$ to be non-envious in the final allocation. There are at most $2^{f(n)}$ such guesses. If an agent $a$ in $K$ shares an edge of $E_{\mathcal{A}}$ with an agent $a^{\prime}$ in $\mathcal{A}_{1}\cup\mathcal{A}_{2}$, then $\phi(a^{\prime})\notin\mathcal{P}_{a}$. To model this, we trim the feasible set $\mathcal{F}_{a^{\prime}}$ to $\mathcal{F}_{a^{\prime}}\setminus\mathcal{P}_{a}$. Call this modified feasibility set $(\mathcal{F}^{\prime}_{a})_{a\in\mathcal{A}_{1}\cup\mathcal{A}_{2}}$. Note that even agents in $R\setminus K$ could turn out to be non-envious, but this does not affect the correctness of the algorithm as some other guess of $K$ will match the exact subset of envious agents in a minimum envy allotment $\phi$.
Let $a^{\prime}$ be an agent in $\mathcal{A}_{1}\cup\mathcal{A}_{2}$. Assume that there is an agent $a\in S$ such that the guessed allocation $\phi(a)$ is in $\mathcal{P}_{a^{\prime}}$. Therefore, $a^{\prime}$ becomes envious if and only if $a^{\prime}$ does not get a house in $\mathcal{P}_{a^{\prime}}$. We can therefore mark $a^{\prime}$ as ‘angry’ and include it in the set $B$, call this new set $B^{\prime}$. This allows us to delete the information about the exact allocation in $S$ and recurse into independent subproblems in $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. To do this, we guess a subset of houses $\mathcal{H}_{1}\subseteq(\mathcal{H}\setminus\phi(S))$ of size $|\mathcal{A}_{1}|$ which are allocated to agents in $\mathcal{A}_{1}$. There are $\binom{n}{|\mathcal{A}_{1}|}$ many such guesses. Then $\mathcal{H}_{2}=\mathcal{H}\setminus(\phi(S)\cup\mathcal{H}_{1})$ is the set of houses allocated to $\mathcal{A}_{2}$. We update $\mathcal{P}_{a}$ to $\mathcal{P}_{a}\cap\mathcal{H}_{1}$ if $a\in\mathcal{A}_{1}$, or to $\mathcal{P}_{a}\cap\mathcal{H}_{2}$ if $a\in\mathcal{A}_{2}$; call this updated preferences as $\mathcal{P}^{\prime}_{a}$. Let us denote by $\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{1}]$ the induced graph of $\mathcal{G}_{\mathcal{A}}$ by the set $\mathcal{A}_{1}$, and similarly for $\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{2}]$.
Let ‘OHAANR’ be a procedure that outputs the minimum envy for the problem of Annnotated Optimal House Allocation of Agent Network. If $\zeta$ is the minimum envy corresponding to the guesses made by the algorithm so far, then we get the following relation:
$\zeta=$ $|D|+|Q|+|R\setminus K|$ $[\text{envious agents in }S]$
$+\text{OHAANR}(\mathcal{A}_{1},\mathcal{H}_{1},\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{1}],(\mathcal{P}^{\prime}_{a})_{a\in\mathcal{A}_{1}},(\mathcal{F}^{\prime}_{a})_{a\in\mathcal{A}_{1}},B^{\prime}\cap\mathcal{A}_{1})$ $[\text{envious agents in }\mathcal{A}_{1}]$
$+\text{OHAANR}(\mathcal{A}_{2},\mathcal{H}_{2},\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{2}],(\mathcal{P}^{\prime}_{a})_{a\in\mathcal{A}_{2}},(\mathcal{F}^{\prime}_{a})_{a\in\mathcal{A}_{2}},B^{\prime}\cap\mathcal{A}_{2})$ $[\text{envious agents in }\mathcal{A}_{2}]$
Moreover, the minimum envy for the problem of Annnotated Optimal House Allocation of Agent Network in the original instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}},(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B)$ is basically the minimum of $\zeta$ over all guesses of $(\phi(a))_{a\in S}$, $\mathcal{H}_{1}$ and $K$.
$\text{OHAANR}(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}},(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B)=\min\limits_{(\phi(a))_{a\in S},\mathcal{H}_{1},K}\big{\{}\zeta\big{\}}$
This allows us to design a recursive algorithm. The correctness of this algorithm is trivial from its design, as it recursively enumerates all possible allocations. We now analyze its running time.
Let $T(n)$ be the maximum running time that this algorithm takes on an input with $n$ agents and $m=n$ houses. Finding a balanced separator of size $f(n)$ takes time at most $\binom{n}{f(n)}\cdot n^{2}$. The total number of choices of $\phi(a)$ for $a\in S$ is at most $n^{f(n)}$. $\mathcal{H}_{1}$ has $\binom{n}{|\mathcal{A}_{1}|}\leqslant 2^{n-1}$ choices. $K$ has $2^{f(n)}$ many choices. All other operations (except the recursion) take time polynomial in $n$. Therefore, they together take time:
$\binom{n}{f(n)}\cdot(2n)^{f(n)}\cdot\binom{n}{|\mathcal{A}_{1}|}\cdot n^{\mathcal{O}(1)}=2^{\mathcal{O}(\log n)\cdot f(n)+(n-1)}$
After fixing $\phi(a)$ for $a\in S$, sets $K$ and $\mathcal{H}_{1}$, the recursed instances take time at most $T(|\mathcal{A}_{1}|)$ and $T(|\mathcal{A}_{2}|)$. Since $|\mathcal{H}_{1}|=|\mathcal{A}_{1}|\leqslant 2n/3$ and $|\mathcal{H}_{2}|=|\mathcal{A}_{2}|\leqslant 2n/3$, we can upper bound both of these expressions by $T(2n/3)$. Moreover, $T(1)=n^{\mathcal{O}(1)}$. This gives us the following recursive relation on the running time.
$T(n)$ $=2^{\mathcal{O}(\log n)\cdot f(n)+(n-1)}\cdot 2T(2n/3)=2^{\mathcal{O}(\log n)\cdot f(n)+n}\cdot T(2n/3)$
$=2^{\mathcal{O}(\log n)\cdot(f(n)+f(2n/3)+f(4n/9)+f(8n/27)+\cdots)+(n+2n/3+4n/9+8n/27+\cdots)}$
$=2^{\mathcal{O}(\log n)\cdot(f(n)+f(2n/3)+f(4n/9)+f(8n/27)+\cdots)+3n}$
For arbitrary $f(n)$, we can bound $(f(n)+f(2n/3)+f(4n/9)+f(8n/27)+\cdots)$ by $f(n)\mathcal{O}(\log n)$. This gives us $T(n)=2^{\mathcal{O}(f(n)\cdot\log^{2}n+n)}$. As mentioned earlier, for $m\neq n$, the running time would become $2^{m+\mathcal{O}(f(n)\cdot\log^{2}n+n)}$.
The pseudocode of this algorithm, named Algorithm 3 is given below.
Algorithm 3 Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$ Output: Minimum envy
1:procedure OHAANR($\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}},(\mathcal{F}_{a})_{a\in\mathcal{A}},B$)
2:     if $\mathcal{A}=\emptyset$ then $\triangleright$ Agent set is empty
3:         return 00 $\triangleright$ No agent is envious
4:     end if
5:     $\xi\leftarrow\infty$triangleright In the worst case, no assignment exists respecting $(\mathcal{F}_{a})_{a\in\mathcal{A}}$
6:     $S\leftarrow$ a balanced separator of size $f(n)$
7:     $\mathcal{A}_{1},\mathcal{A}_{2}\leftarrow$ the separated parts of $\mathcal{A}\setminus S$triangleright $|\mathcal{A}_{1}|,|\mathcal{A}_{2}|\leqslant 2n/3$
8:     for all choices of $\phi(a)\in\mathcal{F}_{a}$ over all $a\in S$ do
9:         $C\leftarrow\{a\in S\mid\phi(a)\in\mathcal{P}_{a}\}$triangleright Agents with preferred house
10:         $D\leftarrow\{a\in S\setminus B\mid\phi(a)\notin\mathcal{P}_{a},\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$triangleright Envious, non-angry
11:         $Q\leftarrow\{a\in S\cap B\mid a\notin\mathcal{P}_{a}\}$triangleright Envious, angry agents
12:         $R\leftarrow S\setminus(C\cup D\cup Q)$triangleright Other agents
13:         $B^{\prime}\leftarrow B\cup\{a^{\prime}\in\mathcal{A}\setminus S\mid\exists a\in S,\{a,a^{\prime}\}\in E_{\mathcal{A}},\phi(a)\in\mathcal{P}_{a^{\prime}}\}$triangleright Updated angry agents
14:         for $\mathcal{H}_{1}\subseteq\mathcal{H}\setminus\phi(S)$ do
15:              $\mathcal{H}_{2}\leftarrow\mathcal{H}\setminus(\phi(S)\cup\mathcal{H}_{1})$
16:              for $K\subseteq R$ do
17:                  for $a\in\mathcal{A}_{1}$ do
18:                       $\mathcal{F}^{\prime}_{a}\leftarrow\mathcal{F}_{a}\cap\mathcal{H}_{1}$
19:                       $\mathcal{P}^{\prime}_{a}\leftarrow\mathcal{P}_{a}\cap\mathcal{H}_{1}$
20:                  end for
21:                  for $a\in\mathcal{A}_{2}$ do
22:                       $\mathcal{F}^{\prime}_{a}\leftarrow\mathcal{F}_{a}\cap\mathcal{H}_{2}$
23:                       $\mathcal{P}^{\prime}_{a}\leftarrow\mathcal{P}_{a}\cap\mathcal{H}_{2}$
24:                  end for
25:                  for $a\in K$, $a^{\prime}\in\mathcal{A}\setminus S$ such that $\{a,a^{\prime}\}\in E_{\mathcal{A}}$ do
26:                       $\mathcal{F}^{\prime}_{a^{\prime}}\leftarrow\mathcal{F}^{\prime}_{a^{\prime}}\setminus\mathcal{P}_{a}$
27:                  end for
28:                  $\xi_{1}\leftarrow\text{OHAANR}(\mathcal{A}_{1},\mathcal{H}_{1},\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{1}],(\mathcal{P}^{\prime}_{a})_{a\in\mathcal{A}_{1}},(\mathcal{F}^{\prime}_{a})_{a\in\mathcal{A}_{1}},B^{\prime}\cap\mathcal{A}_{1})$
29:                  $\xi_{2}\leftarrow\text{OHAANR}(\mathcal{A}_{2},\mathcal{H}_{2},\mathcal{G}_{\mathcal{A}}[\mathcal{A}_{2}],(\mathcal{P}^{\prime}_{a})_{a\in\mathcal{A}_{2}},(\mathcal{F}^{\prime}_{a})_{a\in\mathcal{A}_{2}},B^{\prime}\cap\mathcal{A}_{2})$
30:                  $\xi\leftarrow\min(\xi,|D|+|Q|+|R\setminus K|+\xi_{1}+\xi_{2})$
31:              end for
32:         end for
33:     end for
34:     return $\xi$
35:end procedure
Again, the algorithm can be trivially backtracked to provide the exact allocation $\phi$. Moreover, for all assignments of minimum envy, if we wish to maximize the number of agents getting a preferred house, this could also be done simply by maximizing $C=\{a\in S\mid\phi(a)\in\mathcal{P}_{a}\}$ among all guesses of $C$ that lead to the same envy.
Remark 4.8.
Algorithm 3 can be trivially adapted to output the actual allocation in the same asymptotic runtime. Moreover, the algorithm can also be adapted solve Optimally Happy House Allocation of Agent Network.
[…]
$T(n)$ $=2^{\mathcal{O}\left(\log n\cdot\left(\frac{n}{\log n}+\frac{2n/3}{\log(2n/3)}+\frac{4n/9}{\log(4n/9)}+\cdots\right)+n\right)}$
$=2^{\mathcal{O}\left(\log n\cdot\left(\frac{n}{\log n}+\frac{2n/3}{(3/4)\log(n)}+\frac{4n/9}{(9/16)\log(n)}+\cdots\right)+n\right)}$ as ($\text{as }(3/4)^{r}\log n\leqslant\log((2/3)^{r}n)$
$=2^{\mathcal{O}\left(\log n\cdot\left(\frac{n}{\log n}+\frac{2n/4}{\log n}+\frac{4n/16}{\log n}+\cdots\right)+n\right)}=2^{\mathcal{O}(n)}$
In particular, these include planar graphs, bounded genus graphs [8], and any proper minor closed graph classes [13].
Corollary 4.9.
Consider an input $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ with $m$ houses and $n$ agents and with the guarantee that $m=\mathcal{O}(n)$. Then Algorithm 3 solves Annnotated Optimal House Allocation of Agent Network and hence Optimal House Allocation of Agent Network with the following worst case running times,
[…]
    $2^{\mathcal{O}(n)}$ for $\mathcal{O}\left(\frac{n}{\log n}\right)$-balanced-separable graphs.
5 Complexities for various Vertex Cover sizes of the Agent Graph
Input graphs with small vertex covers are particularly interesting to explore. Such graphs frequently arise in modeling subordinate-supervisor relationships in infrastructures with limited number of supervisors. In this Section, we first design an exact algorithm for Optimal House Allocation of Agent Network and Optimally Happy House Allocation of Agent Network with a running time of $\mathcal{O}((2m)^{k}\cdot(n+m)^{\mathcal{O}(1)})$, where $m$ is the number of houses, $n$ is the number of agents in the agent graph $\mathcal{G}_{\mathcal{A}}$ and $k$ is the size of a vertex cover of $\mathcal{G}_{\mathcal{A}}$. Note that this algorithm performs better than Algorithm 2 when $k=o\left(\frac{n+|E_{\mathcal{A}}|}{\log m}\right)$, where $|E_{\mathcal{A}}|$ is the number of edges in the agent graph. This algorithm also has quasi-polynomial running time when the input agent graphs have polylogarithmic-sized vertex covers. On the other hand, we also show that the problems are $\mathsf{NP}$-hard even when the agent graph is a split graph or a bipartite graph with a vertex cover of size $n^{\varepsilon}$, where $n$ is the number of agents, and $\varepsilon\in(0,1)$ is a constant.
Let the graph on agents $\mathcal{A}$ be $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$. We look into a slice-wise polynomial (XP) algorithm solving Optimal House Allocation of Agent Network with the vertex cover number as the parameter.
Theorem 5.1.
Consider an input $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ with $m$ houses and $n$ agents. There exists an algorithm solving Optimal House Allocation of Agent Network in time $\mathcal{O}((2m)^{k}\cdot(n+m)^{\mathcal{O}(1)})$, where $k$ is the size of a vertex cover of $\mathcal{G}_{\mathcal{A}}$.
Proof 5.2.
Given a set of $n$ agents $\mathcal{A}$, $\mathcal{G}_{\mathcal{A}}=(\mathcal{A},E_{\mathcal{A}})$, set of $m$ houses $\mathcal{H}$, preference $(\mathcal{P}_{a})_{a\in\mathcal{A}}$, assume that a vertex cover $S\subseteq\mathcal{A}$ of $\mathcal{G}_{\mathcal{A}}$ is given, such that $|S|=k$ (if not, this can be computed trivially in $n^{k+\mathcal{O}(1)}$ time). Consider the following algorithm.
[…]
    Guess $\phi(a)$, for all $a\in S$. (there are $\mathcal{O}(m^{k})$ such guesses).
[…]
    Compute $J\subseteq S$, the set of agents already envious due to $\phi(S)$. Precisely,
    $J=\{a\in S\mid\phi(a)\notin\mathcal{P}_{a}\text{ and }\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$
[…]
    Guess happy agents $C\subseteq S\setminus J$ to be the set of agents who will remain non-envious after assigning houses to agents in $\mathcal{A}\setminus S$ (there are $\mathcal{O}(2^{k})$ such guesses).
[…]
    We define a weighted, complete, bipartite graph $\mathcal{G}_{\mathcal{F}}((\mathcal{A}\setminus S)\cup(\mathcal{H}\setminus\phi(S)),E_{\mathcal{F}},w)$ as follows.
    • $(\mathcal{A}\setminus S)$ and $(\mathcal{H}\setminus\phi(S))$ are the two parts. $E_{\mathcal{F}}=\{\{a,h\}\mid a\in(\mathcal{A}\setminus S),h\in(\mathcal{H}\setminus\phi(S))\}$.
    • Set the weights $w(a,h)$ of the edge $\{a,h\}$ to model whether $a$ becomes envious on getting $h$ or not. This can be achieved by doing the following:
        $w(a,h)$ is set to $\infty$, when the assignment of $h\in\mathcal{H}\setminus\phi(S)$ to $a\in\mathcal{A}\setminus S$ is not consistent with the guess of happy agents $C$. This happens when there is some agent $a^{\prime}\in C$, such that $a^{\prime}$ is a neighbour of $a$ but $a^{\prime}$ is envious of $a$. Formally, $w(a,h)=\infty$ when there exists $a^{\prime}\in(C\cap N_{\mathcal{G}_{\mathcal{A}}}(a))$ such that $\phi(a^{\prime})\notin\mathcal{P}_{a^{\prime}}$ but $h\in\mathcal{P}_{a^{\prime}}$.
        $w(a,h)$ is set to $1$, when the assignment of $h$ to $a$ causes $a$ to become envious. Note that since we have already considered the allocation to be consistent with the guess of $C$, no agent in $C$ would be envious of $a$ getting $h$. This happens when there is some agent $a^{\prime}\in S$, such that $a^{\prime}$ is a neighbour of $a$ gets a house preferred by $a$ but $h$ is not preferred by $a$. Formally, $w(a,h)=1$ when there exists $a^{\prime}\in(S\cap N_{\mathcal{G}_{\mathcal{A}}}(a))$ such that $\phi(a^{\prime})\in\mathcal{P}_{a}$ but $h\in\mathcal{P}_{a}$.
        When none of these is the case, the assignment of $h$ to $a$ does not increase envy. We set $w(a,h)=0$.
[…]
    Compute the minimum cost maximum matching (left exhausting) of $\mathcal{G}_{\mathcal{F}}$ and obtain the cost to be $\zeta$. This equates to the total number of envious agents in corresponding assignment of the agents in $\mathcal{A}\setminus S$.
[…]
    For some guess of $C$, the total number of envious agents in $S$ is at most $(k-|C|)$, while the total number of envious agents in $\mathcal{A}\setminus S$ is $\zeta$. Since we are iterating over all guesses of $C$, for any minimum-envy allocation, there would exist a guess $C$ such that $(k-|C|)$ is exactly the number of envious agents. Hence, to solve Optimal House Allocation of Agent Network, it suffices to output the minimum of $k-|C|+\zeta$ over all guesses.
This algorithm takes time $\mathcal{O}((2m)^{k}\cdot(n+m)^{\mathcal{O}(1)})$ and essentially enumerates all possible allocations of $S$ while finding a corresponding minimum-envy extension of that.
We refer to this as Algorithm 4 and provide its pseudocode below.
Algorithm 4 Input: $\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}}$, $S$ Output: Minimum envy
1:$\xi\leftarrow n$triangleright In the worst case, all agents are envious
2:for guess $(\phi(a))_{a\in S}$ do
3:     $J\leftarrow\{a\in S\mid\phi(a)\notin\mathcal{P}_{a}\text{ and }\exists a^{\prime}\in S\cap N_{\mathcal{G}_{\mathcal{A}}}(a),\phi(a^{\prime})\in\mathcal{P}_{a}\}$triangleright already envious
4:     for guess $C\subseteq S\setminus J$ do $\triangleright$ envious in $S$
5:         for ($a\in(\mathcal{A}\setminus S)$, $h\in(\mathcal{H}\setminus\phi(S)$)) do
6:              if there exists $a^{\prime}\in(C\cap N_{\mathcal{G}_{\mathcal{A}}}(a))$ such that $\phi(a^{\prime})\notin\mathcal{P}_{a^{\prime}}$ but $h\in\mathcal{P}_{a^{\prime}}$ then
7:                  $w(a,h)\leftarrow\infty$
8:              else if there exists $a^{\prime}\in(S\cap N_{\mathcal{G}_{\mathcal{A}}}(a))$ such that $\phi(a^{\prime})\in\mathcal{P}_{a}$ but $h\in\mathcal{P}_{a}$ then
9:                  $w(a,h)\leftarrow 1$
10:              else
11:                  $w(a,h)\leftarrow 0$
12:              end if
13:         end for
14:         $\mathcal{G}_{\mathcal{F}}\leftarrow$ complete bipartite graph with parts $(\mathcal{A}\setminus S)$ and $(\mathcal{H}\setminus\phi(S))$ and weights $w$
15:         $\zeta\leftarrow$ minimum cost maximum bipartite matching of $\mathcal{G}_{\mathcal{F}}$
16:         $\xi\leftarrow\min(\xi,k-|C|+\zeta)$
17:     end for
18:end for
19:return $\xi$
In Algorithm 4, if we modify the weight function as
$w^{\prime}(a,h)=\begin{cases}w(a,h)-\frac{1}{n+1}&\text{if }h\notin P_{a}\\ w(a,h)&\text{otherwise}\end{cases}$
we would find a minimum envy allotment that maximizes the number of agents getting a preferred house, solving Optimally Happy House Allocation of Agent Network.
Remark 5.3.
Algorithm 4 could be modified slightly to solve Optimally Happy House Allocation of Agent Network with the same asymptotic running time.
This immediately gives us a polynomial time algorithms for star graphs, as they have vertex cover of size $1$. Moreover this translates to a quasi-polynomial time algorithm for graphs having vertex covers of size $(\log n)^{\mathcal{O}(1)}$; hence we do not expect the problem to be $\mathsf{NP}$-hard for such graphs.
This motivates us to explore the exact boundary of tractable and intractable cases for various sizes of the minimum vertex cover. A natural question would be to ask whether vertex covers of size at most $n^{\varepsilon}$ for any constant $\varepsilon\in(0,1)$ would also allow efficient algorithms. We show in the next couple of theorems, that this is indeed not the case; Optimal House Allocation of Agent Network in $\mathsf{NP}$-hard even for bipartite graphs and split graphs with a vertex cover size of at most $n^{\varepsilon}$ for every constant $\varepsilon\in(0,1)$. Inspired from the reductions due to Madathil et. al. [15], the reductions we provide work even when each house is preferred by constantly many agents.
Theorem 5.4.
Let $\varepsilon\in(0,1)$ be any constant. Consider an input $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ with $m$ houses and $n$ agents. Optimal House Allocation of Agent Network is $\mathsf{NP}$-hard even when $\mathcal{G}_{\mathcal{A}}$ is a complete bipartite graph with one of its partition being of size at most $n^{\varepsilon}$, and each house is preferred by at most four agents.
Proof 5.5.
We reduce from the $\mathsf{NP}$-hard problem of CLIQUE. Let $(G(V,E),k)$ be an instance of CLIQUE, where the problem asks to decide if $G$ has a clique of size $k$. Let $|V|=N$ and $|E|=M$. Moreover, we can assume $N\geqslant(2M)^{\lceil 1/\varepsilon\rceil}$; indeed if that is not the case, we can add (in polynomial time), $(2M)^{\lceil 1/\varepsilon\rceil}$ many isolated vertices to $G$ without changing the size of the maximum clique.
We reduce $(G(V,E),k)$ to an instance of Optimal House Allocation of Agent Network as follows.
[…]
    Let $\mathcal{A}=\{a_{v}\mid v\in V\}\cup\{a^{1}_{e}\mid e\in E\}\cup\{a^{2}_{e}\mid e\in E\}$. That is, we have an agent $a_{v}$ for every vertex $v$ of $G$, and two agents $a^{1}_{e}$, and $a^{2}_{e}$ for each edge $e$ of $G$. Therefore, $n=|\mathcal{A}|=N+2M$.
[…]
    Let =$\mathcal{H}=\{h_{e}\mid e\in E\}\cup\{h^{*}_{j}\mid j\in[N+2M-\binom{k}{2}]\}$. Therefore, $|\mathcal{H}|=M+N+2M-\binom{k}{2}=|\mathcal{A}|+(M-\binom{k}{2})$.
[…]
    We are now going to set the houses $h_{e}$ to be preferred by agents $a^{1}_{e}$ and $a^{2}_{e}$, and also by $a_{u}$ and $a_{v}$ where $e=\{u,v\}$. This is effectively done by setting $\mathcal{P}_{a^{1}_{e}}=\mathcal{P}_{a^{2}_{e}}=\{h_{e}\}$, and $\mathcal{P}_{a_{v}}=\{h_{e^{\prime}}\mid e^{\prime}\in E\text{ and }v\in e^{\prime}\}$. The houses $\{h^{*}_{j}\mid j\in[N+2M-\binom{k}{2}]\}$ are preferred by no agents, i.e. these are dummy houses.
[…]
    We now define the underlying graph $\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}})$. Let $\mathcal{A}_{V}=\{a_{v}\mid v\in V\}$ and let $\mathcal{A}_{E}=\{a^{1}_{e}\mid e\in E\}\cup\{a^{2}_{e}\mid e\in E\}$. Then $\mathcal{G}_{\mathcal{A}}$ defines the complete bipartite graph between $\mathcal{A}_{V}$ and $\mathcal{A}_{E}$. That is, $E_{\mathcal{A}}=\{\{a,a^{\prime}\}\mid a\in\mathcal{A}_{V},a^{\prime}\in\mathcal{A}_{E}\}$.
[…]
    The target number of envious agents is $k$.
Notice that the part $\mathcal{A}_{E}$ of the complete bipartite graph $\mathcal{G}_{\mathcal{A}}$ satisfies $|\mathcal{A}_{E}|=2M\leqslant N^{\varepsilon}=|\mathcal{A}_{V}|^{\varepsilon}\leqslant|\mathcal{A}|^{\varepsilon}$. This reduction can be done in polynomial time. We now show the correctness of the reduction.
Forward direction. Let $(G(V,E),k)$ be a Yes-instance of CLIQUE, i.e., there is clique $S\subseteq V$ of size $k$. Let $E_{S}$ be the edges of the clique $S$, $|E_{S}|=\binom{k}{2}$. Consider the allocation $\phi$ defined as:
[…]
    A dummy house is assigned to $a_{v}$ for all $v\in V$. This uses $N$ dummy houses.
[…]
    A dummy house is assigned to $a^{2}_{e}$ for all $e\in E$. This uses $M$ dummy houses.
[…]
    A dummy house is assigned to $a^{1}_{e}$ for every $e\in E\setminus E_{S}$. This uses $(M-\binom{k}{2})$ dummy houses.
[…]
    For every $e\in E_{S}$,the house $h_{e}$ is assigned to $a^{1}_{e}$. Hence the houses $\{h_{e}\mid e\in E\setminus E_{S}\}$ are all unallocated.
Note that for every $e\in E$, the agents $a^{1}_{e}$ and $a^{2}_{e}$ are all non-envious, because their neighbours in $\mathcal{G}_{\mathcal{A}}$, i.e. the agents $\{a_{v}\mid v\in V\}$ are all assigned dummy houses. Moreover, for all $v\in V$, such that $v\notin S$, no edge in $E_{S}$ is incident to $v$; thus all houses preferred by $a_{v}$ are unallocated, making $a_{v}$ non-envious. Therefore, all envious agents are in $\{a_{v}\mid v\in S\}$. This makes the number of envious agents in $\phi$ to be at most $|S|=k$; the Optimal House Allocation of Agent Network instance is a Yes-instance.
Reverse direction. Let $\phi$ be some allocation with at most $k$ envious agents. We say that $\phi$ is nice if for all $e\in E$, $h_{e}$ is either unassigned or is assigned to $a^{1}_{e}$. If $\phi$ is not nice to begin with, then $\phi$ can be converted to a nice allocation by a sequence of exchanges, none of which increases total envy.
Let $h_{e}$ be some house which is neither unallocated, nor allocated to $a^{1}_{e}$.
[…]
    Case I: $h_{e}$ is assigned to $a^{2}_{e}$. We swap the houses allocated to $a^{1}_{e}$ and $a^{2}_{e}$. This does not change the number of envious agents, due to the symmetry of $\mathcal{G}_{\mathcal{A}}$.
[…]
    Case II: $h_{e}$ is assigned to $a^{j}_{e^{\prime}}$ for some $j\in\{1,2\}$, $e^{\prime}\neq e$. We swap the houses allocated to $a^{1}_{e}$ and $a^{j}_{e^{\prime}}$. Again, this does not increase the number of envious agents. This is because
    1. 1.
      $a^{1}_{e}$ is not envious in both allocations.
    2. 2.
      $a^{j}_{e^{\prime}}$ did not have their preferred house before the swap. Since the neighbourhood of $a^{j}_{e^{\prime}}$ is unaffected by the swap, $a^{j}_{e^{\prime}}$ cannot become envious after the swap, if $a^{j}_{e^{\prime}}$ was not envious before the swap.
    3. 3.
      For all other agents, they are envious after the swap if and only if they were envious before the swap. This is because the set of houses allocated to their neighbours remain unchanged.
[…]
    Case III: $h_{e}$ is assigned to $a_{v}$ for some $v\in V$. Agents $a^{1}_{e},a^{2}_{e}$ must be envious of $a_{v}$ in $\phi$. We now look at the sequence of agents $a^{1}_{e_{0}}=a^{1}_{e},a^{1}_{e_{1}},\ldots,a^{1}_{e_{p-1}},a^{1}_{e_{p}}$ such that
    • $\phi(a_{v})=h_{e_{0}}$.
    • $\phi(a^{1}_{e_{0}})=h_{e_{1}},\phi(a^{1}_{e_{1}})=h_{e_{2}},\ldots,\phi(a^{1}_{e_{p-1}})=h_{e_{p}}$.
    • $\phi(a^{1}_{e_{p}})=h^{*}_{q}$ is a dummy house.
    Note that such a sequence must exist and end with an agent who is assigned a dummy house, as the number of agents and houses are finite. Therefore, $p$ is a finite integer. Now, we modify the allocation to $\phi^{\prime}$ as follows:
    • $\phi^{\prime}(a_{v})=h^{*}_{q}$.
    • $\phi^{\prime}(a^{1}_{e_{0}})=h_{e_{0}},\phi^{\prime}(a^{1}_{e_{1}})=h_{e_{1}},\ldots,\phi^{\prime}(a^{1}_{e_{p-1}})=h_{e_{p-1}},\phi^{\prime}(a^{1}_{e_{p}})=h_{e_{p}}$.
    • $\phi^{\prime}(a)=\phi(a)$ for all other agents $a$.
    We will now argue why this set of exchanges do not increase the number of envious agents.
    • Case A: $e$ is incident on $v$. Let $e=\{u,v\}$. In $\phi$, $a_{v}$ was not envious, but $a_{u}$ may or may not be envious. The agents $a^{1}_{e}$, and $a^{2}_{e}$ were both envious. In $\phi^{\prime}$, $a^{1}_{e}$ and $a^{2}_{e}$ are not envious, agent $a_{v}$ becomes envious, and $a_{u}$ may become envious. Therefore, the number of envious agents in $\{a^{1}_{e},a^{2}_{e},a_{u},a_{v}\}$ do not increase. The agents $a^{1}_{e_{1}},a^{1}_{e_{2}},\ldots,a^{1}_{e_{p}}$ all become non-envious in $\phi^{\prime}$ as they get their preferred houses. For all other agents $a^{j}_{e^{\prime}}$ where $e^{\prime}$ is an edge in $E$, the dummy house $h^{*}_{q}$ is the only house which appears in the set of houses assigned to neighbours of $a^{j}_{e^{\prime}}$ in $\phi^{\prime}$ which did not appear in $\phi$; hence $a^{j}_{e^{\prime}}$ is envious in $\phi^{\prime}$ only if they were envious in $\phi$. Finally, all other agents $a_{v^{\prime}}$ where $v^{\prime}$ is a vertex in $V$, are envious in $\phi^{\prime}$ if and only if they are envious in $\phi$ as only the house $h_{v}$ (a house not preferred by $a_{v^{\prime}}$) gets added to the set of houses assigned to the neighbours of $a_{v^{\prime}}$.
    • Case B: $e$ is not incident on $v$. Let $e=\{u_{1},u_{2}\}$. In $\phi$, the agents $a_{u_{1}}$, $a_{u_{2}}$ may or may not be envious; but the agents $a^{1}_{e}$ and $a^{2}_{e}$ were both envious. In $\phi^{\prime}$, $a^{1}_{e}$ and $a^{2}_{e}$ are not envious, while agents $a_{u_{1}}$ and $a_{u_{2}}$ may become envious. Therefore, the number of envious agents in $\{a_{u_{1}},a_{u_{2}},a^{1}_{e},a^{2}_{e}\}$ do not increase. The agents $a^{1}_{e_{1}},a^{1}_{e_{2}},\ldots,a^{1}_{e_{p}}$ all become non-envious in $\phi^{\prime}$ as they get their preferred houses. If the agent $a_{v}$ is not envious in $\phi$, then it cannot become envious in $\phi^{\prime}$ as only $h_{e}$ (a house not preferred by $a_{v}$) gets added to the set of houses assigned to its neighbours. For all other agents $a^{j}_{e^{\prime}}$ where $e^{\prime}$ is an edge in $E$, the dummy house $h^{*}_{q}$ is the only house which appears in the set of houses assigned to neighbours of $a^{j}_{e^{\prime}}$ in $\phi^{\prime}$ which did not appear in $\phi$; hence $a^{j}_{e^{\prime}}$ is envious in $\phi^{\prime}$ only if they were envious in $\phi$. Finally, all other agents $a_{v^{\prime}}$ where $v^{\prime}$ is a vertex in $V$, are envious in $\phi^{\prime}$ if and only if they were envious in $\phi$ as only the house $h_{v}$ (a house not preferred by $a_{v^{\prime}}$) gets added to the set of houses assigned to the neighbours of $a_{v^{\prime}}$.
These exchanges do not increase the total envy, but they allocate at least one house $h_{e}$ to $a^{1}_{e}$, which was previously allocated to some other agent. Repeating this for at most $|\mathcal{H}|$ times would give us a nice allocation, with envy no more than that in $\phi$.
Hence, we can safely assume that $\phi$ is a nice allocation with envy at most $k$. $\phi$ allocates dummy houses to $a_{v}$ for all $v\in V$ and to $a^{2}_{e}$ for all $e\in E$. Moreover, it is safe to assume that all dummy houses are assigned to some agent, otherwise we can assign a dummy house to any $a^{1}_{e}$ without increasing total envy.
We now look into agents which are not assigned dummy vertices by $\phi$. There are $N+2M-(N+2M-\binom{k}{2})=\binom{k}{2}$ many such agents, each of the form $a^{1}_{e}$ for some $e\in E$. Let $\gamma=\binom{k}{2}$. Let $a^{1}_{e_{1}},a^{1}_{e_{2}},\ldots,a^{1}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. For any $v\in V$, $v$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. Further, the number of envious agents, i.e. the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore indeed, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.
This completes the proof of correctness of the reduction.
Next we explore split graphs. Split graphs are graphs that can be partitioned into a clique and an independent set. Note that the clique can act as a vertex cover. We show the the problem Optimal House Allocation of Agent Network, even when restricted to split graphs of $n$-vertices having the clique of size at most $n^{\varepsilon}$, for any constant $\varepsilon\in(0,1)$, is $\mathsf{NP}$-hard.
Theorem 5.6.
Let $\varepsilon\in(0,1)$ be any constant. Consider an input $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}}),(\mathcal{P}_{a})_{a\in\mathcal{A}})$ with $m$ houses and $n$ agents. Optimal House Allocation of Agent Network is $\mathsf{NP}$-hard even when $\mathcal{G}_{\mathcal{A}}$ is a split graph where the maximum clique is of size at most $n^{\varepsilon}$, and each house is preferred by at most three agents.
Proof 5.7.
We provide a reduction from the CLIQUE problem to Optimal House Allocation of Agent Network. Let $(G(V,E),k)$ be an instance of CLIQUE, which asks to decide whether $G$ has a clique of size $k$.
Let $|V|=N,|E|=M\geqslant 1$. Define $t=N^{\lceil 1/\varepsilon\rceil}$. We now construct an instance of Optimal House Allocation of Agent Network as follows:
[…]
    Let $\mathcal{A}=\{a_{v}\mid v\in V\}\cup\{a^{j}_{e}\mid e\in E,j\in[t]\}$. That is, we have an agent $a_{v}$ for every vertex $v$ of $G$, and $t$ agents $a^{1}_{e},a^{2}_{e},\ldots,a^{t}_{e}$ for each edge $e$ of $G$. Therefore, $n=|\mathcal{A}|=N+M\cdot t=N+M\cdot N^{\lceil 1/\varepsilon\rceil}$.
[…]
    Let =$\mathcal{H}=\{h^{j}_{e}\mid e\in E,j\in[t]\}\cup\{h^{*}_{j}\mid j\in[(M-\binom{k}{2})t+N]\}$. Therefore, $|\mathcal{H}|=M\cdot t+N+(M-\binom{k}{2})t=|\mathcal{A}|+(M-\binom{k}{2})t$.
[…]
    We are now going to set the houses $h^{j}_{e}$ to be preferred by agents $a^{j}_{e}$, $a_{u},a_{v}$, where $e=\{u,v\}$. This is effectively done by setting $\mathcal{P}_{a^{j}_{e}}=\{h^{j}_{e}\}$, and $\mathcal{P}_{a_{v}}=\{h^{j}_{e}\mid j\in[t],e\in E\text{ such that }v\in e\}$. The houses $\{h^{*}_{j}\mid j\in[(M-\binom{k}{2})t+N]\}$ are preferred by no agents, i.e., these are dummy houses.
[…]
    We now define the underlying graph $\mathcal{G}_{\mathcal{A}}(\mathcal{A},E_{\mathcal{A}})$. The edges in $E_{\mathcal{A}}$ are defined as follows:
    • For every $u,v\in V$, $\{a_{u},a_{v}\}\in E_{\mathcal{A}}$.
    • For every $v\in V,e\in E,j\in[t]$, $\{a_{v},a^{j}_{e}\}\in E_{\mathcal{A}}$.
[…]
    The target number of envious agents is $k$.
This means that $C=\{a_{v}\mid v\in V\}$ forms a clique and $I=\{a^{j}_{e}\mid e\in E,j\in[t]\}$ forms an independent set with $\mathcal{A}=C\cup I$. Moreover, $|C|=N<(N+M\cdot N^{\lceil 1/\varepsilon\rceil})^{\varepsilon}=n^{\varepsilon}$. The output of the reduction is the Optimal House Allocation of Agent Network instance $(\mathcal{A},\mathcal{H},\mathcal{G}_{\mathcal{A}},(\mathcal{P}_{a})_{a\in\mathcal{A}},k)$. This reduction can be done in polynomial time as $\varepsilon$ is a constant. We now show the correctness of this reduction.
Forward direction. Let $(G(V,E),k)$ be a Yes-instance of CLIQUE, i.e., there is a clique $S\subseteq V$ of size $k$. Let $E_{S}$ be the edges of the clique $S$. We have $|E_{S}|=\binom{k}{2}$. Consider the allocation $\phi$ defined as:
[…]
    A dummy house is assigned to $a_{v}$ for all $v\in V$. This uses $N$ dummy houses.
[…]
    A dummy house is assigned to $a^{j}_{e}$ for every $j\in[t],e\in E\setminus E_{S}$. This uses $(M-\binom{k}{2})t$ dummy houses.
[…]
    For every $j\in[t],e\in E_{S}$, the house $h^{j}_{e}$ is assigned to $a^{j}_{e}$. Hence the houses $\{h^{j}_{e}\mid j\in[t],e\in E\setminus E_{S}\}$ are all unallocated.
Note that for every $e\in E$, the agents $a^{j}_{e}$ are all non-envious, because their neighbours in $\mathcal{G}_{\mathcal{A}}$, i.e., the agents $\{a_{v}\mid v\in V\}$ are all assigned dummy houses. Moreover, for all $v\in V$, such that $v\notin S$, no edge in $E_{S}$ is incident to $v$; thus, all houses preferred by $a_{v}$ are unallocated, making $a_{v}$ non-envious. Therefore, all envious agents are in $\{a_{v}\mid v\in S\}$. This makes the number of envious agents in $\phi$ to be at most $|S|=k$; the Optimal House Allocation of Agent Network instance is a Yes-instance.
Reverse direction. Let $\phi$ be some allocation with at most $k$ envious agents. We say that $\phi$ is nice if for all $j\in[t],e\in E$, $h^{j}_{e}$ is either unassigned or is assigned to $a^{j}_{e}$. If $\phi$ is not nice to begin with, then $\phi$ can be converted to a nice allocation by a sequence of exchanges, none of which increases total envy.
Let $h^{j}_{e}$ be some house which is neither unallocated, nor allocated to $a^{j}_{e}$.
[…]
    Case I: $h^{j}_{e}$ is allocated to $a^{j^{\prime}}_{e^{\prime}}$ for some $j^{\prime}\neq j$ or $e^{\prime}\neq e$. Then we swap houses allocated to $a^{j}_{e}$ and $a^{j^{\prime}}_{e^{\prime}}$. After this swap, $a^{j}_{e}$ is non-envious by getting a preferred house. The agent $a^{j^{\prime}}_{e^{\prime}}$ cannot become envious after the swap if $a^{j^{\prime}}_{e^{\prime}}$ was not envious before the swap; this is because the houses allocated to the neighbours of $a^{j^{\prime}}_{e^{\prime}}$ remain unchanged and $a^{j^{\prime}}_{e^{\prime}}$ did not have a preferred house before the swap. For all other agents, they remain envious after the swap if and only if they were envious before the swap (as the set of houses allocated to their neighbours is unchanged).
[…]
    Case II: $h^{j}_{e}$ is allocated to $a_{v}$ for some $v\in V$. $a^{j}_{e}$ must be envious of $a_{v}$ in $\phi$. We now look at the sequence of agents $a^{j}_{e}=a^{j_{0}}_{e_{0}},a^{j_{1}}_{e_{1}},\ldots,a^{j_{p-1}}_{e_{p-1}},a^{j_{p}}_{e_{p}}$ such that
    • $\phi(a_{v})=h^{j_{0}}_{e_{0}}$.
    • $\phi(a^{j_{0}}_{e_{0}})=h^{j_{1}}_{e_{1}},\phi(a^{j_{1}}_{e_{1}})=h^{j_{2}}_{e_{2}},\ldots,\phi(a^{j_{p-1}}_{e_{p-1}})=h^{j_{p}}_{e_{p}}$.
    • $\phi(a^{j_{p}}_{e_{p}})=h^{*}_{q}$ is a dummy house.
    Note that such a sequence must exist and end with an agent who is assigned a dummy house, as the number of agents and houses are finite. Therefore, $p$ is a finite integer. Now, we modify the allocation to $\phi^{\prime}$ as follows:
    • $\phi^{\prime}(a_{v})=h^{*}_{q}$.
    • $\phi^{\prime}(a^{j_{0}}_{e_{0}})=h^{j_{0}}_{e_{0}},\phi^{\prime}(a^{j_{1}}_{e_{1}})=h^{j_{1}}_{e_{1}},\ldots,\phi^{\prime}(a^{j_{p-1}}_{e_{p-1}})=h^{j_{p-1}}_{e_{p-1}},\phi^{\prime}(a^{j_{p}}_{e_{p}})=h^{j_{p}}_{e_{p}}$.
    • $\phi^{\prime}(a)=\phi(a)$ for all other agents $a$.
    In $\phi^{\prime}$, $a^{j}_{e}$ is not envious anymore, however, $a_{v}$ may or may not become envious. The agents $a^{j_{1}}_{e_{1}},\ldots,a^{j_{p-1}}_{e_{p-1}},a^{j_{p}}_{e_{p}}$, were allocated their preferred house in $\phi^{\prime}$, and hence are not envious. Moreover, for all other $a^{j^{\prime}}_{e^{\prime}}$, they are envious in $\phi^{\prime}$ if and only if they were envious in $\phi$; this is because the dummy house $h^{*}_{q}$ is the only house that appears in the set of houses allocated to their neighbours in $\phi^{\prime}$ but not in $\phi$. The same happens for $a_{v^{\prime}}$, $v^{\prime}\in V\setminus\{v\}$; the set of houses allocated to the neighbours of $a_{v^{\prime}}$ does not change. Therefore this does not increase total envy.
These exchanges do not increase the total envy, but they allocate at least one house $h^{j}_{e}$ to $a^{j}_{e}$, which was previously allocated to some other agent. Repeating this for at most $|\mathcal{H}|$ times would give us a nice allocation, with envy no more than that in $\phi$.
Hence, we can safely assume that $\phi$ is a nice allocation with envy at most $k$. $\phi$ allocates dummy houses to $a_{v}$ for all $v\in V$. Moreover, it is safe to assume that all dummy houses are assigned to some agent, otherwise we can assign a dummy house to any $a^{j}_{e}$ without increasing total envy.
We now look into agents which are not assigned dummy vertices by $\phi$. There are $Mt-(M-\binom{k}{2})t=\binom{k}{2}t$ many such agents. Let $\gamma=\binom{k}{2}t$. Let $a^{j_{1}}_{e_{1}},a^{j_{2}}_{e_{2}},\ldots,a^{j_{\gamma}}_{e_{\gamma}}$ be the agents which are not assigned dummy houses. Therefore, the number of distinct edges in $e_{1},e_{2},\ldots,e_{\gamma}$ is at least $\gamma/t=\binom{k}{2}$. Moreover, for any $v\in V$, $v$ is an endpoint of at least one edge in $e_{1},e_{2},\ldots,e_{\gamma}$, if and only if the agent $a_{v}$ is envious. Therefore, the number of envious agents is precisely the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on. The number of vertices that $\binom{k}{2}$ distinct edges are incident on, is at least $k$, with equality holding if and only if the edges induce a clique of size $k$. However, the number of envious agents, i.e., the number of vertices on which the edges $e_{1},e_{2},\ldots,e_{\gamma}$ are incident on, is at most $k$ by assumption. Therefore, $\{e_{1},e_{2},\ldots,e_{\gamma}\}$ is a set of $\binom{k}{2}$ distinct edges, which induce a clique of size $k$.
This completes the proof of correctness of the reduction.
6 Conclusion
We study the Optimal House Allocation of Agent Network and Optimally Happy House Allocation of Agent Network problem, natural extensions of the classical House Allocation problem to social networks.
Our results reveal a rich complexity theoretic landscape. We present polynomial-time algorithms when each agent has a unique preferred house. Then, we show NP-hardness results under structural restrictions, including split graphs, complete bipartite graphs, and 3-regular graphs with uniform preferences. We further look into structural properties of the agent graphs and design exact (exponential) algorithms exploiting these structures, that are more efficient than the naive brute-force algorithm.
An interesting problem for future work would be to analyze the following decision problem: Given a set of agents $\mathcal{A}$, a set of houses $\mathcal{H}$, an underlying agent graph $\mathcal{G}_{\mathcal{A}}$, preferences $(\mathcal{P}_{a})_{a\in\mathcal{A}}$, and integers $k$, $b$, determine if there is an allocation that achieves envy at most $k$ and happiness at least $b$. Note that this is at least as hard as Optimally Happy House Allocation of Agent Network.